### v1

# Stackelberg vs. Nash in the Lottery Colonel Blotto Game

Abstract
Resource competition problems are often modeled using Colonel Blotto games. However, Colonel Blotto games only simulate scenarios where players act simultaneously. In many real-life scenarios, competition is sequential, such as cybersecurity, cloud services, business investments, and more. To model such sequential competition, we model the Lottery Colonel Blotto game as a Stackelberg game. We solve the Stackelberg equilibrium in the Lottery Colonel Blotto game in which the first mover’s strategy is actually a solution to a bi-level optimization problem. We develop a constructive method that allows for a series of game reductions. This method enables us to compute the leader’s optimal commitment strategy in a polynomial number of iterations. Furthermore, we identify the conditions under which the Stackelberg equilibrium aligns with the Nash equilibria. Finally, we show that by making the optimal first move, the leader’s utility can be unbounded compared to its utility in Nash equilibria. We find that the player with a smaller budget has a greater incentive to become the leader in the game. Surprisingly, even when the leader adopts the optimal commitment strategy, the follower’s utility may improve compared to that in Nash equilibria.
1 Introduction
Competing for resources with limited budgets in strategic settings has a wide range of applications. For example:
Advertising auctions and e-Commerce markets. Advertisers compete for limited ad placements on online platforms, allocating their budgets to maximize exposure and conversion rates. Large enterprises often act as leaders, with smaller retailers responding as followers. These budget strategies significantly impact the ad visibility and conversions.
Economic aspects of Web, GenAI, and cloud computing. Providers like AWS and Google Cloud compete across multiple markets (e.g., geographic regions or industries), allocating cloud resources to attract customers and maximize market share. One provider acts as the leader, allocating resources first to focus on high-value markets. The other, as the follower, observes the leader’s strategy and optimizes its own allocation, targeting underdeveloped or less saturated markets.
These examples, along with many others from areas such as electoral competition (Behnezhad et al., 2018), security (Pagey et al., 2023; Zhao et al., 2023; Ferdowsi et al., 2017; Gupta et al., 2014), crowdsourcing (Roitero et al., 2022; Haggiag et al., 2022; AlShehry and Ferguson, 2015), and recommendation systems (Anwar et al., 2024; Lee et al., 2022), are modeled by Colonel Blotto games (Borel, 1921). These types of games have attracted considerable interest from sociology (Xiao et al., 2020; Fu and Iyer, 2019), economics (Jackson and Nei, 2015), and computer science (Kohli et al., 2012).
There are many variants of Colonel Blotto games, in which players simultaneously allocate their limited budgets (resources) across various battlefields. These budgets can be represented by integers, such as troops, or by real numbers, such as bid prices in auctions. The player’s utility on a single battlefield is determined by the resources invested by players in that battlefield and the outcome function of that battlefield. A player’s payoff is then calculated by aggregating the utility gained from all battlefields. Based on the outcome function of the battlefield, Colonel Blotto games can be categorized into two types: one where the outcome function follows the winner-takes-all rule, and the other where the outcome function is proportional. Colonel Blotto games with a proportional outcome function are also referred to as the Lottery Colonel Blotto game. In this paper, we focus on the Lottery Colonel Blotto game.
Although Colonel Blotto games are highly versatile, most existing research treats them as normal-form games, primarily focusing on equilibrium existence, computation, and closed-form characterizations in restricted settings (Roberson, 2006; Hortala-Vallve and Llorente-Saguer, 2012; Macdonell and Mastronardi, 2015; Perchet et al., 2022). However, in many real-world situations, such as the Web, cloud computing, cybersecurity, and advertising auction examples, one player can commit to a strategy before the other player makes their decision.
In this paper, we model the Lottery Colonel Blotto game as a Stackelberg game. The Stackelberg game describes a strategic scenario in which one player, as the leader (first mover), determines strategy, while the other player, as the follower (second mover), responds based on the leader’s strategy (Hicks, 1935; Schelling, 1960).
The Stackelberg game can be viewed as a sequential game. In many previous studies, the focus has been on scenarios involving an attacker, as a follower, and a defender, as a leader, both operating under cost constraints rather than budget limitations. Under these cost constraints, they determine their respective efforts in order to maximize their individual payoffs (Zhuang and Bier, 2007; Cavusoglu et al., 2008; Hausken et al., 2008; Hausken, 2012). However, Iliaev et al. (2023) point out that the model without budget constraints is easier to solve analytically because it is equivalent to solving a single-battlefield model with $n$ different values. Iliaev et al. (2023) consider budget constraints to compare the sequential and the simultaneous game. However, they only impose budget constraints on the follower, while the first-mover still has only costs without budget limitations, and both sides have symmetric valuations for the same battlefield. Here, for broader applicability, we consider a more general model in which both the leader and the follower have budgets. In addition, both players and the battlefield values are asymmetric. Furthermore, we will compare the sequential game and the simultaneous game, i.e., the Stackelberg equilibrium vs. Nash equilibria.
In addition to the above dynamics, there exists another type of dynamics. All players simultaneously allocate resources in the first battlefield, then allocate resources in the second battlefield, and so on (Xie and Zheng, 2022; Anbarci et al., 2023).
To compare sequential games and simultaneous games, we need to analyze Stackelberg equilibrium and Nash equilibria. However, finding a closed-form representation of Nash equilibria in Colonel Blotto games is a known challenge (Perchet et al., 2022; Li and Zheng, 2022). Such representations have been characterized only in specific scenarios, such as when battlefield values are identical (Roberson, 2006) or when players have symmetrical budgets (Hortala-Vallve and Llorente-Saguer, 2012; Boix-Adserà et al., 2020). Most existing research offers only partial characterizations, focusing on conditions for the existence and uniqueness of Nash equilibria. For the Stackelberg equilibrium, we need to compute the leader’s optimal commitment. However, previous work has provided negative conclusions. Finding an optimal pure strategy to commit to is NP-hard in the normal-form games (Conitzer and Sandholm, 2006; Korzhyk et al., 2010; Letchford and Conitzer, 2010), as it requires solving a bi-level optimization problem (Renou, 2009). To address this, we investigate new insights into optimal commitment and best-response dynamics, enabling a comparison between Stackelberg equilibrium and Nash equilibria in these games.
1.1 Our Contribution
Our contributions are multi-fold:
  • We offer a novel understanding of the follower’s best response using a water-filling approach.
  • We construct a series of game reductions by splitting battlefields, ensuring that the follower’s valuation for each sub-battlefield is uniform. We have proven that there exists an injective mapping between players’ strategies before and after each reduction of the game. Additionally, the players’ utilities remain unchanged. By analyzing the leader’s commitment in the reduced game, we characterize the optimal commitment in the original game. Ultimately, we reduce the support of the follower’s best response strategy from $12^{n}-$ 1 to $n$, where $n$ is the number of battlefields.
  • By reformulating the leader’s objective and reducing the search space for the follower’s best response strategy, we can characterize the leader’s optimal commitment strategy through a polynomial number of iterations.
  • We characterize the sufficient and necessary conditions, under which the Stackelberg equilibrium coincides with the Nash equilibria of the simultaneous-move game.
  • Finally, we address our motivating questions by constructing extreme cases that compare the leader’s utility when using the optimal commitment strategy to their utility in Nash equilibria. We observe that this ratio is also dependent on the leader’s budget relative to the follower’s budget. Furthermore, we provide an example that illustrates how the follower’s utility may still increase when faced with the leader’s optimal commitment.
1.2 Related Work
Colonel Blotto Games and Their Variants. There is an extensive body of literature on Colonel Blotto games.
When the outcome function is winner-takes-all, the player who allocates the most resources to a battlefield wins that battlefield. In this line of work, a pure Nash equilibrium does not always exist, while a mixed Nash equilibrium always does (Roberson, 2006; Macdonell and Mastronardi, 2015). Specifically, each player’s strategy is a complex joint distribution over an n-dimensional simplex. Roberson (2006) provides a closed-form solution for the Nash equilibrium in the case of two players and multiple battlefields with equal value. Macdonell and Mastronardi (2015) provide a detailed and comprehensive analysis of the Nash equilibrium in the case of two players and two battlefields with different values. In view of the difficulty in analyzing Nash equilibrium, researchers propose the General Lotto game, a well-known variant. Specifically, each player’s strategy is a joint distribution over the n-dimensional simplex, ensuring that the expected allocation of resources does not exceed the budget, rather than strictly staying within the budgets (Hart, 2008; Dziubiński, 2013; Kovenock and Roberson, 2021). The Nash equilibrium solution of the General Lotto game has been obtained (Kovenock and Roberson, 2021).
When the outcome function is winner-takes-all, the probability of a player winning a battlefield depends on the proportion of resources they allocate compared to the total resources allocated by all players. This type of Colonel Blotto games is also called the Lottery Colonel Blotto game. In this line of work, it is called symmetric if all players have equal budgets; otherwise, it is asymmetric. When all battlefields have the same value, and this value is consistent for all players, the game is homogeneous. If the battlefields possess different values, but these values remain the same for every player, the game is heterogeneous. Finally, when the battlefields’ values differ, and these values vary for each player, the game is termed a generalized game. The Lottery Colonel Blotto game has been proven to possess pure Nash equilibria (Kim et al., 2018). Friedman (1958) shows that the pure Nash equilibrium is unique and exhibits proportionality features in a two-player setup with heterogeneous battlefield values and asymmetric player budgets. Duffy and Matros (2015) extend Friedman’s analysis from two players to multiple players. Kim et al. (2018) provide a method to identify all pure Nash equilibria in the case of two players with asymmetric budgets and multiple battlefields with generalized values, and present an example where the pure Nash equilibrium is not unique. Kovenock and Arjona (2019) provide the best response given the strategy of the other player. However, none of the existing work has provided a closed-form representation of the pure Nash equilibria (Kim et al., 2018; Kovenock and Arjona, 2019). In a variant where the outcome function of a battlefield is determined by the Tullock rent-seeking contest success function, the proportional rule is parameterized by a battle-specific discriminatory power and a battle-and-contestant-specific lobbying effectiveness (Xu and Zhou, 2018; Osorio, 2013; Li and Zheng, 2022). In this variant, the existence of a pure Nash equilibrium is not always guaranteed. Xu et al. (2022) provide sufficient conditions for the existence of the pure Nash equilibria, while Li and Zheng (2022) analyze the properties of the pure Nash equilibria and provide sufficient conditions for its uniqueness.
Sequential Colonel Blotto Games. There are two types of sequentiality: one is the sequentiality of the actions taken by players, and the other is the sequentiality of the appearance of the battlefields.
In the sequentiality of the actions taken by players, the first player makes a move, followed by the second player. Many studies have compared sequential games and simultaneous games, in other contexts. For instance, Zhuang and Bier (2007) identify equilibrium strategies for both attacker and defender in simultaneous and sequential games, although their model does not consider players’ budgets. Chandan et al. (2020) consider three-stage Colonel Blotto games with stronger and weaker players. In their model, the weaker player has the option to pre-commit resources to a single battlefield of its choice, and the stronger player can choose whether to allocate resources to win that battlefield. Chandan et al. (2022) propose a two-stage General Lotto game, in which one of the players has reserved resources that can be strategically pre-allocated across the battlefields in the first stage and the players then engage by simultaneously allocating their real-time resources against each other.
In the sequentiality of the appearance of the battlefields, all players simultaneously allocate resources in the first battlefield, then allocate resources in the second battlefield, and so on. Anbarci et al. (2023) analyze dynamic Blotto contests, where battlefields are presented to players in a predetermined sequential order. They focus on the sub-game perfect equilibrium, exploring the existence and uniqueness of this solution concept. Xie and Zheng (2022) construct a pure strategy Markov perfect equilibrium (when it exists) and provide closed-form solutions for players’ strategies and winning probabilities. Klumpp et al. (2019) explore the strategic allocation of resources in a dynamic setting where winning a majority of battlefields is the goal. They provide the optimal strategies for both players in sub-game perfect equilibrium. Li and Zheng (2021) reveal that the even-split strategy is robust when players have incomplete information about the other player’s resource allocation.
Other Variants and Applications. The Colonel Blotto games and their variants, along with the Tullock contest, all-pay auction, and their generalizations, can find wide application in various domains. These include competition design (Deng et al., 2023), contest design (Ghosh and Kleinberg, 2014; Levy et al., 2017; Letina et al., 2023; Dasgupta and Nti, 1998), and auctions (Tang et al., 2016; Brânzei et al., 2012).
2 Preliminaries
In the Stackelberg model of the Lottery Colonel Blotto game, let $a$ and $b$ represent the leader and the follower, respectively. Both players have limited budget constraints, where $x_{a}>0$ and $x_{b}>0$. These players allocate their budgets across $n$ battlefields, represented by the set $[n]=\{1,2,\dots,n\}$. For player $i\in\{a,b\}$, let $x_{ij}\geq 0$ denote the budget invested by player $i$ in battlefield $j$. Throughout this paper, we consider players’ pure strategies, which are denoted as $\bm{x}_{i}=(x_{i1},x_{i2},\cdots,x_{in})$. The strategy set for player $i$ is: $\bm{X}_{i}\triangleq\{\bm{x}_{i}:\sum_{j=1}^{n}x_{ij}=x_{i}\text{ and }x_{ij}\geq 0\}.$
Player $i\in\{a,b\}$ assigns a value to battlefield $j$ as $v_{ij}\in\mathbb{Q}_{>0}$. Hence, the game can be represented as
$\mathcal{G}:=\langle\{a,b\},[n],x_{a},x_{b},(v_{aj})_{j=1}^{n},(v_{bj})_{j=1}^{n}\rangle.$
Given a strategy profile $(\bm{x}_{a},\bm{x}_{b})$, the utility of player $i\in\{a,b\}$ on battlefield $j\in[n]$ is defined as the proportion of the budget allocated by player $i$ relative to the total budgets allocated by both players. That is,
$u_{ij}(x_{ij},x_{-ij})=\frac{x_{ij}}{x_{ij}+x_{-ij}}\cdot v_{ij}.$ (1)
In the event that both players allocate zero budget to a battlefield $j$, we assume that the follower will win the entire battlefield. This assumption is based on the rationale that the follower can achieve this outcome by allocating even an arbitrarily small amount of budget to $j\in[n]$. This assumption guarantees the existence of the follower’s best response strategy. Using a linear aggregation function, player $i$’s utility in the game is given by:
$u_{i}(\bm{x}_{a},\bm{x}_{b})=\sum_{j=1}^{n}u_{ij}=\sum_{j=1}^{n}\frac{x_{ij}}{x_{ij}+x_{-ij}}\cdot v_{ij},\quad\forall i\in\{a,b\}.$ (2)
When $x_{aj}>0$, for all $j\in[n]$, Kovenock and Arjona (2019) characterize the other player’s best response functions as follows. This characterization is a useful tool for our subsequent analysis. To use this result to characterize the leader’s optimal commitment, we first assume that the leader will allocate a positive budget to every battlefield. Following a series of constructions in Section 3, we identify the support of the follower’s best response strategies and provide a closed-form expression of the leader’s optimal commitment strategy in Section 4. To complete this analysis, we will verify that the leader indeed allocates a positive budget to every battlefield in the optimal commitment strategy. We focus on the pure strategy of the leader.
Lemma 1.
(Kovenock and Arjona, 2019) Given the leader’s strategy $\bm{x}_{a}=(x_{aj})_{j=1}^{n}$, where $x_{aj}>0$, $\forall j\in[n]$, assume without loss of generality that all battlefields are ordered such that $\frac{v_{b1}}{x_{a1}}\geq\frac{v_{b2}}{x_{a2}}\geq\cdots\geq\frac{v_{bn}}{x_{an}}$. The unique optimal budget allocation of the follower, $\bm{x}_{b}=(x_{b1},x_{b2},\cdots,x_{bn})$, to battlefield $j\in[n]$ is characterized as follows:
$x_{bj}=\begin{cases}\frac{\left(x_{aj}v_{bj}\right)^{\frac{1}{2}}\cdot\left(x_{b}+\sum_{j^{\prime}\in K(\bm{x}_{a})}x_{aj^{\prime}}\right)}{\sum_{j^{\prime}\in K(\bm{x}_{a})}\left(x_{aj^{\prime}}v_{bj^{\prime}}\right)^{\frac{1}{2}}}-x_{aj},&\text{ if }j\in K(\bm{x}_{a});\\ 0,&\text{ if }j\in[n]\backslash K(\bm{x}_{a}),\end{cases}$ (3)
where $K(\bm{x}_{a})=\{1,\cdots,k^{*}\}$ is such that
$k^{*}=\max\left\{k\in[n]:\frac{v_{bj}}{x_{aj}}>\frac{\left(\sum_{l=1}^{j}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l=1}^{j}x_{al}\right)^{2}},\quad\forall j\leq k\right\}.$ (4)
The uniqueness of the follower’s best response $\bm{x}_{b}$ will facilitate the subsequent analysis. However, providing a closed-form representation of the leader’s optimal commitment strategy remains very challenging. This difficulty primarily arises from the uncertainty with regard to which battlefields the follower will abandon, specifically where $x_{bj}=0$ for $j\in[n]\backslash K(\bm{x}_{a})$. For simplicity, we denote $\overline{K(\bm{x}_{a})}=[n]\backslash K(\bm{x}_{a})$ and the support $k^{*}=|K(\bm{x}_{a})|$. In the following section, we characterize the support of the follower’s best response strategy when the leader employs optimal commitment strategy. This analysis, in turn, will aid in fully computing the leader’s optimal commitment strategy.
Due to space limitations, most proofs are deferred to the Appendix. However, we provide technical remarks and insights into the proofs within the main text.
3 The Support of the Follower’s Best Response Strategy
The computation of the leader’s optimal commitment strategy can be formulated as a bi-level optimization problem. In this scenario, the leader, assuming the follower is a utility maximizer, selects an optimal strategy in the upper-level optimization task, while the follower provides the best response in the lower-level optimization task. The complexity is compounded by the fact that the support of each player’s strategy profile can have up to $12^{n}-$ 1 possible combinations. Additionally, it is challenging to determine the exact budget allocation for each battlefield.
In this section, we demonstrate that when the leader employs optimal commitment strategy, the support of the follower’s best response strategy is limited to at most $n$ possible combinations. This significantly reduces the search space for the leader’s optimal commitment strategy.
We achieve this characterization by constructing a series of auxiliary games.
  1. 1.
    Consider a game $\mathcal{G}$, with the leader’s commitment strategy denoted by $\bm{x}_{a}$ (which may not be optimal), and the follower’s best response denoted by $\bm{x}_{b}$. If we divide a single battlefield into multiple sub-battlefields in such a way that the leader and follower’s valuations of the original battlefield are evenly distributed among these sub-battlefields, the resulting game is denoted by $\mathcal{G}^{(1)}$. By evenly distributing their budgets across these sub-battlefields, the utilities for both the leader and the follower remain unchanged in $\mathcal{G}^{(1)}$. The formal statement and its proof are deferred to Lemma 7 in the Appendix A due to space constraints.
  2. 2.
    By repeatedly performing this battlefield-splitting process, we create a game in which the follower values all battlefields equally. Without loss of generality, we can rename the battlefields so that the leader values them in increasing order. Denote this game by $\mathcal{G}^{(2)}$. According to Lemma 8 in the Appendix A, in the optimal commitment strategy, the leader should allocate more or at least equal budget to battlefields that they value strictly higher in order to achieve a strictly greater payoff.
  3. 3.
    Furthermore, we observe that, in the optimal commitment strategy, when the leader values two battlefields equally, they should allocate an equal budget to both battlefields. For a formal statement and proof, see Lemma 9 in the Appendix A.
  4. 4.
    Following the above procedure, we obtain a game where $v_{b1}=v_{b2}=\dots$, meaning the follower values all battlefields equally. Meanwhile, the leader values the battlefields in increasing order, with some values strictly increasing and others remaining equal. Our final operation is to merge these battlefields that split from the same battlefield. In this new game $\mathcal{G}^{(3)}$, the leader and follower’s values and budgets across these sub-battlefields are merged into the original battlefield. Their utilities remain unchanged from $\mathcal{G}^{(2)}$. The formal statement and proof are deferred to Lemma 10 in the Appendix A.
With these constructions and observations, we can now prove the main result of this section.
Theorem 1.
In a Stackelberg game
$\mathcal{G}:=\langle\{a,b\},[n],x_{a},x_{b},(v_{aj})_{j=1}^{n},(v_{bj})_{j=1}^{n}\rangle,$
assume without loss of generality that the battlefields are ordered such that $\frac{v_{a1}}{v_{b1}}\leq\frac{v_{a2}}{v_{b2}}\leq\cdots\leq\frac{v_{an}}{v_{bn}}$. Then, when the leader employs the optimal commitment strategy $\bm{x}_{a}$, the support of the follower’s best response strategy has at most $n$ possibilities. Specifically, $K(\bm{x}_{a})\in\{\{1\},\{1,2\},\dots,\{1,2,\dots,n\}\}$. Additionally, if for $j,h\in[n]$, $\frac{v_{aj}}{v_{bj}}=\frac{v_{ah}}{v_{bh}}$, then the optimal commitment $\bm{x}_{a}$ satisfies $\frac{x_{aj}}{x_{ah}}=\frac{v_{aj}}{v_{ah}}$.
Technical Remark: The constructions in this section to derive the main result rely on an interpretation different from the one provided by Kovenock and Arjona (2019). To prove Lemma 1, they represent the second player’s best response strategy as the solution to an optimization problem. By verifying that (3) satisfies the Kuhn-Tucker conditions, which are both necessary and sufficient, they conclude that (3) is the unique global constrained maximizer of the problem.
In contrast, we interpret the follower’s optimization problem as a water-filling process. Initially, the follower allocates their budget to the battlefields with the highest marginal utility. Note that the first and second derivatives of the follower’s utility from battlefield $j$ are:
$\frac{\partial u_{bj}}{\partial x_{bj}}=\frac{\partial(\frac{x_{bj}v_{bj}}{x_{aj}+x_{bj}})}{\partial x_{bj}}=\frac{x_{aj}v_{bj}}{(x_{aj}+x_{bj})^{2}},\quad\frac{\partial^{2}u_{bj}}{\partial x_{bj}^{2}}=\frac{-2x_{aj}v_{bj}}{(x_{aj}+x_{bj})^{3}}<0.$ (5)
Hence, the utility $u_{bj}$ is a concave function with respect to $x_{bj}$. As the budget allocation to battlefield $j$ increases, its marginal utility decreases until it matches the marginal utility of other battlefields with initially lower marginal utility. From this point, the follower distributes their budget across these battlefields, maintaining equal marginal utility, until it eventually decreases to an even lower level. Ultimately, the follower exhausts their entire budget. The lowest marginal utility is given by $\frac{\left(\sum_{l=1}^{k^{*}}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l=1}^{k^{*}}x_{al}\right)^{2}}$, as defined in $k^{*}$ in (4).
The proofs of Lemmas 7 through 10 involve applying the follower’s best response strategy. It is essential to consider whether these game-splitting and merging operations, as well as the leader’s strategies $\bm{x}_{a}$, alter the structure of the set $K(\bm{x}_{a})$. This set, $K(\bm{x}_{a})$, represents the battlefields on which the follower allocates positive budgets. Essentially, there is a mapping between the leader’s commitment set and the follower’s strategy set $\bm{x}_{b}$ across the original game $\mathcal{G}$ and the subsequent games $\mathcal{G}^{(k)}$ for $k=1,2,3$.
4 The Leader’s Optimal Commitment Strategy
Theorem 1 characterizes $K(\bm{x}_{a})$, the support of the follower’s best response strategy when the leader employs their optimal commitment strategy $\bm{x}_{a}$. In this section, we formulate the leader’s objective as an optimization problem for any given set $K(\bm{x}_{a})$. By solving the optimization problem for all $n$ possible realizations of the set $K(\bm{x}_{a})$ as described in Theorem 1, and comparing the optimal values, we identify the optimal commitment strategy $\bm{x}_{a}$.
Assume the support of the follower’s best response strategy, when the leader employs their optimal commitment strategy, is given by $K$. Given a specific set $K(\bm{x}_{a})$, the leader’s optimal commitment strategy can be determined by solving the following optimization problem (OC).
[…]
s.t. $\sum_{j=1}^{n}x_{aj}=x_{a},$ (7)
$x_{aj}>0,\forall j\in[n],$ (8)
$x_{bj}=\frac{(x_{aj}\cdot v_{bj})^{\frac{1}{2}}\cdot(\sum_{j^{\prime}\in K}x_{aj^{\prime}}+x_{b})}{\sum_{j^{\prime}\in K}(x_{aj^{\prime}}\cdot v_{bj^{\prime}})^{\frac{1}{2}}}-x_{aj}>0,\forall j\in K,$ (9)
$x_{bj}=0,\forall j\in\overline{K}.$ (10)
The leader’s utility is derived from two components: the battlefield set $K$, where the leader and the follower share the battlefield proportional to their resources allocated, and the set $\overline{K}$, where the leader has completely won the battlefield. Equation (7) represents the leader’s budget constraint. Inequalities (8) assume the leader allocates a positive amount of budget to every battlefield in the optimal commitment strategy. Constraints (9) and (10) describe the follower’s best response strategy as outlined in Lemma 1.
Next, we characterize the budget that the leader allocates to the battlefields where the follower does not compete. This characterization is based on interpreting the follower’s best response strategy through a water-filling approach, but from the leader’s perspective. Consider any battlefield $j\in\overline{K}$. On one hand, the leader must allocate sufficient budget to these battlefields so that the follower’s marginal utility on these battlefields is lower than on other battlefields, making competition unappealing. On the other hand, the leader does not need to allocate an excessive amount of resources to force the follower to withdraw. Therefore, there exists a minimum threshold of resources that the leader must allocate to these battlefields in $\overline{K}$, which is sufficient but not wasteful. This threshold value is provided below.
Lemma 2.
In the leader’s optimal commitment $\bm{x}_{a}$, we have $x_{aj}=\frac{v_{bj}\cdot\left(x_{b}+\sum_{l\in K(\bm{x}_{a})}x_{al}\right)^{2}}{\left(\sum_{l\in K(\bm{x}_{a})}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},\forall j\in\overline{K(\bm{x}_{a})}.$
Lemma 1 provides us with $\frac{v_{bj}}{x_{aj}}\leq\frac{\left(\sum_{l\in K(\bm{x}_{a})}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l\in K(\bm{x}_{a})}x_{al}\right)^{2}}$, $\forall j\in\overline{K(\bm{x}_{a})}$. This implies $x_{aj}\geq\frac{v_{bj}\cdot\left(x_{b}+\sum_{l\in K(\bm{x}_{a})}x_{al}\right)^{2}}{\left(\sum_{l\in K(\bm{x}_{a})}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}$. To demonstrate that equality holds for any $j\in\overline{K(\bm{x}_{a})}$, we compare the leader’s utility before and after shifting an arbitrarily small amount of budget $\epsilon$ from a battlefield $j\in\overline{K(\bm{x}_{a})}$ to a battlefield $j\in K(\bm{x}_{a})$. If this budget shift consistently increases the leader’s utility, the leader will allocate precisely the amount stated in the lemma to keep the set $\overline{K(\bm{x}_{a})}$ unchanged. The full proof of the lemma is deferred to the Appendix B.
Then, we simplify the optimization problem (OC) in three aspects: (i) we add the expression for $x_{aj},j\in\overline{K}$ as a constraint in (OC); (ii) we substitute $x_{bj}$ as outlined in (9) and (10) into the objective function (6), thereby removing $x_{bj}$ from the leader’s optimization problem; (iii) we note that the second term of (6) can be removed, as it becomes a constant when the set $\overline{K}$ is fixed. Denote the objective $\hat{u}_{a}(\bm{x}_{a})=u_{a}(\bm{x}_{a},\bm{x}_{b})-\sum_{j\in\overline{K}}v_{aj}$.
Therefore, given the set $K(\bm{x}_{a})$, the optimization problem (OC) can be reformulated as follows.
[…]
s.t. $x_{aj}=\frac{v_{bj}\cdot\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}\cdot v_{bl})^{\frac{1}{2}}\right)^{2}},\quad\forall j\in\overline{K},$ (12)
$\sum_{j=1}^{n}x_{aj}=x_{a},$ (13)
$x_{aj}>0,\quad\forall j\in[n].$
Solving (OC’) is still challenging, as it is not a typical convex programming problem that can be solved in polynomial time. Additionally, heuristics that approximate the optimal solution do not help us in ultimately comparing the Stackelberg equilibrium with the Nash equilibria. To address this challenge, we develop the following characterization of the leader’s optimal commitment strategy $x_{aj}$, when $j\in K$.
Lemma 3.
Let $\bm{x}_{a}$ be the leader’s optimal commitment. There exist two parameters, $\alpha$ and $\beta$, such that $\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}}=\sqrt{x_{aj}}\beta$, for all $j\in K(\bm{x}_{a})$.
Together with (12) and (13), we can eliminate the parameter $\beta$ in Lemma 3 by establishing an equation through the leader’s optimal commitment strategy $\bm{x}_{a}$. By substituting $x_{aj}$’s into (11), we reformulate the leader’s utility as a function of the parameter $\alpha$. This transforms the non-convex optimization problem (OC’) into a single-variable function maximization problem. Note that $K(\bm{x}_{a})$ can be one of the sets $\{\{1\},\{1,2\},\dots,\{1,2,\dots,n\}\}$. By considering one possible realization of $K(\bm{x}_{a})$ at a time and taking the derivative of the objective function with respect to $\alpha$, we can identify $x_{aj}$’s as below. We refer to each of these strategies $\bm{x}_{a}$ that corresponds to a set $K(\bm{x}_{a})$ as a candidate optimal commitment strategy.
For the sake of simplicity, we define the following expressions.
Definition 1.
For $i\in\{a,b\}$, $K\subseteq[n]$, let
$v_{iK}$ $=\sum_{j\in K}v_{ij},$
$c_{K}$ $=\sum_{j\in K}\frac{v_{aj}^{2}}{v_{bj}},$
$B_{1}(K)$ $=x_{a}v_{bK}^{2}-2x_{b}v_{bK}v_{b\overline{K}},$
$B_{2}(K)$ $=4x_{b}v_{aK}v_{b\overline{K}}-2x_{a}v_{aK}v_{bK},$
$B_{3}(K)$ $=x_{a}v_{aK}^{2}-2x_{b}v_{b\overline{K}}c_{K},$
$B_{4}(K)$ $=x_{a}^{2}v_{bK}^{2}-4x_{b}(x_{a}+x_{b})v_{bK}v_{b\overline{K}},$
$B_{5}(K)$ $=8x_{b}(x_{a}+x_{b})v_{aK}v_{b\overline{K}}-2x_{a}^{2}v_{aK}v_{bK},$
$B_{6}(K)$ $=x_{a}^{2}v_{aK}^{2}-4x_{b}(x_{a}+x_{b})c_{K}v_{b\overline{K}},$
$\phi_{1}(\theta,K)$ $=B_{1}(K)\theta^{2}+B_{2}(K)\theta+B_{3}(K),$
$\phi_{2}(\theta,K)$ $=B_{4}(K)\theta^{2}+B_{5}(K)\theta+B_{6}(K).$
Theorem 2.
In a Stackelberg game
$\mathcal{G}:=\langle\{a,b\},[n],x_{a},x_{b},(v_{aj})_{j=1}^{n},(v_{bj})_{j=1}^{n}\rangle,$
assume, without loss of generality, that the battlefields are ordered such that $\frac{v_{a1}}{v_{b1}}\leq\frac{v_{a2}}{v_{b2}}\leq\cdots\leq\frac{v_{an}}{v_{bn}}$. For each set $K\in\{\{1\},\{1,2\},\dots,\{1,2,\dots,n\}\}$, solving the optimization problem (OC’) can yield the leader’s candidate optimal commitment strategy $\bm{x}_{a}$ as follows.
Case 1. $\forall j,h\in K$, $\frac{v_{aj}}{v_{bj}}=\frac{v_{ah}}{v_{bh}}$, or $K=\{1\}$, then
$x_{aj}=\begin{cases}x_{aK}\cdot\frac{v_{aj}}{v_{aK}},&\text{if}\;j\in K;\\ \frac{v_{bj}\cdot\left(x_{b}+x_{aK}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},&\text{if}\;j\in\overline{K},\end{cases}$
where $x_{aK}=\frac{x_{a}v_{bK}-2x_{b}v_{b\overline{K}}+\sqrt{x_{a}^{2}v_{bK}^{2}-4x_{a}x_{b}v_{bK}v_{b\overline{K}}-4x_{b}^{2}v_{bK}v_{b\overline{K}}}}{2(v_{b\overline{K}}+v_{bK})}$.
Case 2.1. there exist $j,h\in K$, such that $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, and $K=[n]$, then
$x_{aj}=\frac{\left(\frac{v_{aj}}{\sqrt{v_{bj}}}+(\frac{c_{K}}{v_{bK}})^{\frac{1}{2}}\cdot\sqrt{v_{bj}}\right)^{2}}{\sum_{l=1}^{n}\left(\frac{v_{al}}{\sqrt{v_{bl}}}+(\frac{c_{K}}{v_{bK}})^{\frac{1}{2}}\cdot\sqrt{v_{bl}}\right)^{2}}\cdot x_{a},\quad\forall j\in[n].$
Case 2.2. there exist $j,h\in K$, such that $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, and $K\subsetneq[n]$, then
$x_{aj}=\begin{cases}\frac{2x_{b}^{2}v_{b\overline{K}}\cdot(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{\phi_{1}(\alpha,K)-(v_{aK}-v_{bK}\alpha)\cdot\sqrt{\phi_{2}(\alpha,K)}},\quad&j\in K;\\ \frac{v_{bj}\cdot\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},\quad&j\in\overline{K},\end{cases}$
where $\alpha$ is the optimal solution to the following optimization problem
$\max_{\alpha}\quad$ $\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)\cdot v_{aj}\right)\cdot\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{\frac{\phi_{1}(\alpha,K)-(v_{aK}-v_{bK}\alpha)\cdot\sqrt{\phi_{2}(\alpha,K)}}{2x_{b}^{2}v_{b\overline{K}}}\cdot x_{b}+\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$
s.t. $x_{a}^{2}-4\cdot\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\right)\cdot v_{b\overline{K}}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}\cdot x_{b}(x_{a}+x_{b})\geq 0.$
Proof.
Let $\bm{x}_{a}$ denote the leader’s optimal commitment and $\bm{x}_{b}$ denote the follower’s best response. First, we consider Case 1. Let $x_{aK}=\sum_{l\in K}x_{al}$. By Theorem 1, it is easy to obtain that for $\forall j\in K$, $x_{aj}=x_{aK}\frac{v_{aj}}{v_{aK}}$. For $\forall j\in\overline{K}$, according to Lemma 2, we have
$x_{aj}=\frac{v_{bj}(x_{b}+x_{aK})^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}=\frac{v_{bj}(x_{b}+x_{aK})^{2}}{\left(\sum_{l\in K}(x_{aK}\frac{v_{al}}{v_{aK}}v_{bl})^{\frac{1}{2}}\right)^{2}}=\frac{v_{bj}(x_{b}+x_{aK})^{2}}{x_{aK}v_{bK}}.$
The amount of resources allocated by the leader in the set $\overline{K}$ is given by $\sum_{j\in\overline{K}}x_{aj}=\sum_{j\in\overline{K}}\frac{v_{bj}(x_{b}+x_{aK})^{2}}{x_{aK}v_{bK}}$. Then, we have
$x_{aK}+\sum_{j\in\overline{K}}\frac{v_{bj}(x_{b}+x_{aK})^{2}}{x_{aK}v_{bK}}=x_{a}.$ (14)
Rearranging Equation (14), we obtain
$(v_{b\overline{K}}+v_{bK})(x_{aK})^{2}+(2x_{b}v_{b\overline{K}}-x_{a}v_{bK})x_{aK}+x_{b}^{2}v_{b\overline{K}}=0.$
This is a quadratic equation. It has two positive real roots for $x_{aK}$. In this case, we choose the larger root to maximize the leader’s utility:
$x_{aK}=\frac{x_{a}v_{bK}-2x_{b}v_{b\overline{K}}+\sqrt{x_{a}^{2}v_{bK}^{2}-4x_{a}x_{b}v_{bK}v_{b\overline{K}}-4x_{b}^{2}v_{bK}v_{b\overline{K}}}}{2(v_{b\overline{K}}+v_{bK})}.$
Obviously, it requires that $x_{a}^{2}v_{bK}^{2}-4x_{a}x_{b}v_{bK}v_{b\overline{K}}-4x_{b}^{2}v_{bK}v_{b\overline{K}}\geq 0$. If $x_{a}^{2}v_{bK}^{2}-4x_{a}x_{b}v_{bK}v_{b\overline{K}}-4x_{b}^{2}v_{bK}v_{b\overline{K}}<0$, it means that for the given set $K$, the leader does not have sufficient resources to cover the battlefield sets $\overline{K}$ or $K$. In this case, we should seek the next $K$. Therefore, if $x_{a}^{2}v_{bK}^{2}-4x_{a}x_{b}v_{bK}v_{b\overline{K}}-4x_{b}^{2}v_{bK}v_{b\overline{K}}\geq 0$, the optimal commitment is as follows
$x_{aj}=\begin{cases}x_{aK}\cdot\frac{v_{aj}}{v_{aK}},&\text{if}\;j\in K;\\ \frac{v_{bj}\left(x_{b}+x_{aK}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},&\text{if}\;j\in\overline{K}.\end{cases}$
Second, we consider the case where $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, leading to subcases 2.1 and 2.2.
In this case, $\beta\neq 0$. According to Lemma 3, for $\forall j\in K$, we have $(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\beta}=x_{aj}^{\frac{1}{2}}$. The resources invested by the leader in $K$ are $\frac{1}{\beta^{2}}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}$. For $\forall j\in\overline{K}$, we have
$x_{aj}$ $=v_{bj}\cdot\frac{(x_{b}+x_{K})^{2}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\frac{1}{\beta}\right)^{2}},$
where $x_{K}=\sum_{j\in K}x_{aj}=\frac{1}{\beta^{2}}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})^{2}$. The resources invested by the leader in $\overline{K}$ are
$\sum_{j\in\overline{K}}x_{aj}=\sum_{j\in\overline{K}}v_{bj}\cdot\beta^{2}\frac{(x_{b}+x_{K})^{2}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}.$
We will analyze different values of $K$ and leave this proof to the Appendix B. ∎
The Leader’s Optimal Commitment Strategy. In a Stackelberg game $\mathcal{G}:=\langle\{a,b\},[n],x_{a}$, $x_{b}$, $(v_{aj})_{j=1}^{n},(v_{bj})_{j=1}^{n}\rangle$, following Theorem 2, we can compute the candidate optimal commitment strategy $\bm{x}_{a}$ for each set $K(\bm{x}_{a})\in\{\{1\},\{1,2\},\dots,\{1,2,\dots,n\}\}$. For each candidate strategy $\bm{x}_{a}$, we then determine the leader’s utility. The strategy $\bm{x}_{a}$ that yields the highest utility is identified as the leader’s optimal commitment strategy.
Recall that the best response function provided by Kovenock and Arjona (2019) requires $x_{aj}>0$, for all $j\in[n]$. To conclude this section, we demonstrate that if there exists a battlefield $j$ such that $x_{aj}=0$ in the leader’s optimal commitment strategy, as identified through our process, then increasing $x_{aj}$ to an arbitrarily small budget and carefully adjusting resource allocation in other battlefields will enhance the leader’s utility. Therefore, the leader’s optimal commitment strategy must satisfy this prerequisite.
Lemma 4.
In the leader’s optimal commitment $\bm{x}_{a}$, for $\forall j\in[n]$, $x_{aj}>0$.
5 Stackelberg vs. Nash
In this section, we investigate the necessary and sufficient conditions under which the Stackelberg equilibrium coincides with the Nash equilibria.
Let’s define $\frac{v_{aj}}{v_{bj}}$ as the relative value ratio of the two players over battlefield $j$. Our first observation is that the relative value ratio can only have two distinct values if a Stackelberg equilibrium is also a Nash equilibrium.
Lemma 5.
Let $\bm{x}_{a}$ be the leader’s optimal commitment strategy and $\bm{x}_{b}$ be the follower’s best response to $\bm{x}_{a}$. If $\bm{x}_{a}$ is also the best response strategy to $\bm{x}_{b}$, then the cardinality of the set $\{\frac{v_{aj}}{v_{bj}},\forall j\in[n]\}$ is at most two.
To prove this lemma, we apply Lemma 1 twice, since $\bm{x}_{a}$ and $\bm{x}_{b}$ are best response strategies to each other. Together with Lemma 3, we can establish a quadratic equation whose single variable is $\frac{v_{aj}}{v_{bj}},\forall j\in[n]$. Since a quadratic equation can have at most two distinct roots, the lemma is proved.
Lemma 5 allows us to merge battlefields with identical relative value ratios. This step is essential for proving the main result of this section, as detailed below. First, for simplicity, let’s define a few expressions.
Definition 2.
For any $T\subseteq[n]$, $i\in\{a,b\}$, let
$v_{iT}$ $=\sum_{j\in T}v_{ij},$
$\Psi_{1}(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ $=\left(\xi_{1}\xi_{3}^{-\frac{1}{2}}+\left(\frac{\xi_{1}^{2}\xi_{3}^{-1}+\xi_{2}^{2}\xi_{4}^{-1}}{\xi_{3}+\xi_{4}}\right)^{\frac{1}{2}}\cdot\xi_{3}^{\frac{1}{2}}\right)^{2},$
$t_{1}(T)$ $=\Psi_{1}(v_{aT},v_{a\overline{T}},v_{bT},v_{b\overline{T}}),$
$t_{2}(T)$ $=\Psi_{1}(v_{a\overline{T}},v_{aT},v_{b\overline{T}},v_{bT}),$
$\Psi_{2}(\xi_{1},\xi_{2},\xi_{3},\xi_{4},\xi_{5},\xi_{6})$ $=\frac{\xi_{6}\xi_{4}\xi_{1}(\xi_{5}\xi_{3})^{\frac{1}{2}}}{(\xi_{5}\xi_{3})^{\frac{1}{2}}+(\xi_{6}\xi_{4})^{\frac{1}{2}}}-\frac{\xi_{5}\xi_{3}\xi_{2}(\xi_{6}\xi_{4})^{\frac{1}{2}}}{(\xi_{5}\xi_{3})^{\frac{1}{2}}+(\xi_{6}\xi_{4})\frac{1}{2}},$
$\Psi_{3}(\xi_{1},\xi_{2},\xi_{3},\xi_{4},\xi_{5},\xi_{6})$ $=\frac{\xi_{1}\xi_{4}\xi_{6}\xi_{5}}{\xi_{5}+\xi_{6}}-\frac{\xi_{6}\xi_{5}\xi_{3}\xi_{2}}{\xi_{5}+\xi_{6}},$
$t_{3}(T)$ $=\Psi_{2}(v_{a\overline{T}},v_{aT},v_{b\overline{T}},v_{bT},t_{2}(T),t_{1}(T))$
$\quad+\Psi_{3}(v_{aT},v_{a\overline{T}},v_{bT},v_{b\overline{T}},t_{1}(T),t_{2}(T)),$
$t_{4}(T)$ $=\Psi_{2}(v_{aT},v_{a\overline{T}},v_{bT},v_{b\overline{T}},t_{1}(T),t_{2}(T)),$
$R(T)$ $=\frac{t_{4}(T)}{t_{3}(T)}.$
Theorem 3.
A Stackelberg equilibrium is also a Nash equilibrium if and only if one of the following conditions hold:
  • The relative value ratio is consistent across all battlefields. That is, there exists a constant $c$ such that $\frac{v_{aj}}{v_{bj}}=c$, for $\forall j\in[n]$.
  • There exists a set $M$, such that
    (1) $\frac{v_{aj}}{v_{bj}}=\frac{v_{ak}}{v_{bk}}$ for any $j,k\in M$, and $\frac{v_{aj}}{v_{bj}}=\frac{v_{ak}}{v_{bk}}$ for any $j,k\in\overline{M}=[n]\backslash M$, and (2) $\frac{x_{a}}{x_{b}}=R(T)$.
Note that the leader can always employ the Nash Equilibrium strategy in the Stackelberg game to ensure that their payoff is at least equal to that in the normal form game. Theorem 3 characterizes the scenarios under which being the first mover does not provide any additional benefit.
6 Leader’s Advantage and Impact
In this section, we analyze the advantages of the leader as a first mover, with these advantages being contingent upon the leader’s budget relative to that of the follower. Intuitively, when the leader makes an optimal commitment, the leader’s utility increases, while the follower’s utility tends to decrease. However, we are surprised to find that the follower’s utility can still improve in response to the leader’s optimal commitment.
Player’s strategy in the Nash equilibrium. Li and Zheng (2022) present an approach to compute the Nash equilibrium in two-player Lottery Colonel Blotto game. We can utilize this approach to compute Nash equilibria.
Lemma 6.
[Li and Zheng (2022)] In a two-player Lottery Colonel Blotto game, assume that $\frac{v_{b1}}{v_{a1}}\leq\frac{v_{b2}}{v_{a2}}\leq\cdots\leq\frac{v_{bn}}{v_{an}}$. Let
$f(\mu)=\sum_{h=1}^{n}\left[v_{ah}\left(\frac{v_{bh}}{v_{ah}}\right)\mu\left(\mu-\frac{v_{bh}}{v_{ah}}\frac{x_{a}}{x_{b}}\right)\cdot\prod_{j\neq h}\left(\mu+\frac{v_{bj}}{v_{aj}}\right)^{2}\right],$
and $\mu^{*}$ is a positive solution to $f(\mu)=0$. Then, $(\bm{x^{*}}_{a},\bm{x^{*}}_{b})$ is a Nash equilibrium of the game if and only if
$x_{ah}^{*}$ $=\mu^{*}\cdot\frac{v_{bh}\mu^{*}/\left(\mu^{*}+\left(\frac{v_{bh}}{v_{ah}}\right)\right)^{2}}{\sum_{j=1}^{n}v_{bj}\left(\frac{v_{bj}}{v_{aj}}\right)\mu^{*}/\left(\mu^{*}+\left(\frac{v_{bj}}{v_{aj}}\right)\right)^{2}}\cdot x_{b},$
$x_{bh}^{*}$ $=\frac{v_{bh}\left(\frac{v_{bh}}{v_{ah}}\right)\mu^{*}/\left(\mu^{*}+\left(\frac{v_{bh}}{v_{ah}}\right)\right)^{2}}{\sum_{j=1}^{n}v_{bj}\left(\frac{v_{bj}}{v_{aj}}\right)\mu^{*}/\left(\mu^{*}+\left(\frac{v_{bj}}{v_{aj}}\right)\right)^{2}}\cdot x_{b},$
and $\mu^{*}\in[\min_{h}\frac{v_{bh}}{v_{ah}}\frac{x_{a}}{x_{b}},\max_{h}\frac{v_{bh}}{v_{ah}}\frac{x_{a}}{x_{b}}]$.
The leader’s budget is relatively large. The following theorem gives a lower bound for the leader’s utility in Nash equilibria.
Theorem 4.
For any game
$\mathcal{G}:=\langle\{a,b\},[n],x_{a},x_{b},(v_{aj})_{j=1}^{n},(v_{bj})_{j=1}^{n}\rangle,$
the leader’s utility in the Nash equilibrium is greater than or equal to $\frac{x_{a}}{x_{a}+x_{b}}\cdot(\sum_{j=1}^{n}v_{aj})$.
Proof.
Let $(\bm{x}^{*}_{a},\bm{x}^{*}_{b})$ denote the Nash equilibrium. Consider a leader’s strategy $\hat{\bm{x}}_{aj}=\frac{x_{a}}{x_{b}}x^{*}_{bj}$, we have $u_{a}(\bm{x}^{*}_{a},\bm{x}^{*}_{b})\geq u_{a}(\hat{\bm{x}}_{a},\bm{x}^{*}_{b})$. Furthermore, we have
$u_{a}(\hat{\bm{x}}_{a},\bm{x}^{*}_{b})=\sum_{j=1}^{n}\frac{\frac{x_{a}}{x_{b}}x^{*}_{bj}}{\frac{x_{a}}{x_{b}}x^{*}_{bj}+x^{*}_{bj}}\cdot v_{aj}=\frac{x_{a}}{x_{a}+x_{b}}\sum_{j=1}^{n}v_{aj}.$
It immediately shows that the utility of the leader in the Nash equilibrium is greater than or equal to $\frac{x_{a}}{x_{a}+x_{b}}\cdot(\sum_{j=1}^{n}v_{aj})$. ∎
Corollary 1.
Let $u_{a}^{NE}$ and $u_{a}^{SE}$ denote the leader’s utility under the Nash equilibrium and Stackelberg equilibrium, respectively. We have $\frac{u_{a}^{SE}}{u_{a}^{NE}}\leq\frac{x_{a}+x_{b}}{x_{a}}$
Proof.
Let $(\tilde{\bm{x}}_{a},\tilde{\bm{x}}_{b})$ denote the Stackelberg equilibrium, then by Theorem 4, we have
$\frac{x_{a}}{x_{a}+x_{b}}\sum_{j=1}^{n}v_{aj}\leq u_{a}(\bm{x}^{*}_{a},\bm{x}^{*}_{b})\leq u_{a}(\tilde{\bm{x}}_{a},\tilde{\bm{x}}_{b})<\sum_{j=1}^{n}v_{aj}.$
Rearranging the inequality yields $\frac{u_{a}(\tilde{\bm{x}}_{a},\tilde{\bm{x}}_{b})}{u_{a}(\bm{x}^{*}_{a},\bm{x}^{*}_{b})}\leq\frac{x_{a}+x_{b}}{x_{a}}.$
This corollary shows that when $x_{a}$ is relatively large, $\frac{u_{a}(\tilde{\bm{x}}_{a},\tilde{\bm{x}}_{b})}{u_{a}(\bm{x}^{*}_{a},\bm{x}^{*}_{b})}$ approaches 1. Hence, when the leader’s budget is relatively large, the leader can make a commitment to enhance their own benefits, but the improvement is marginal.
The ratio of the leader’s utility under the Stackelberg equilibrium to that under the Nash equilibrium. We consider a game with two battlefields. Analyzing this two-battlefield scenario can provide insights into the leader’s advantages in games with multiple battlefields. Let $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},x_{a},1,(v_{a1},1),(v_{b1},1)\rangle$. Without loss of generality, let $v_{a1}\leq v_{b1}$. The following theorem gives the range of the ratio.
Theorem 5.
In game $\tilde{\mathcal{G}}$, let $\hat{\bm{x}}_{a}$ denote the leader’s optimal commitment, $\hat{\bm{x}}_{b}$ denote the follower’s best response, and $u_{a}^{NE}$ denote the leader’s utility in the Nash equilibrium. Then, the ratio of Player $a$’s utility in the Stackelberg equilibrium to their utility in the Nash equilibrium is bounded by
$\frac{v_{a1}+1}{v_{a1}+\frac{v_{b1}(x_{a}+1)}{v_{b1}x_{a}+v_{a1}}}\leq\frac{u_{a}(\hat{\bm{x}}_{a},\hat{\bm{x}}_{b})}{u_{a}^{NE}}\leq\frac{v_{a1}+1}{\frac{x_{a}v_{a1}^{2}}{x_{a}v_{a1}+v_{b1}}+\frac{x_{a}}{x_{a}+1}}.$
The leader’s budget is relatively small. Consider an extreme case $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},\epsilon,1,(\epsilon,1),(o(\epsilon),1)\rangle$. We observe that when the leader’s budget is dominated by the follower’s budget, and the follower’s value on the first battlefield is negligible compared to the leader’s value on the first battlefield, the ratio of the leader’s utility when moving first to their utility in the Nash equilibrium approaches infinity. This suggests that the player with weak budget has a stronger incentive to act as the first mover.
The changes in the follower’s utility. When the leader makes an optimal commitment, it is natural to think that the follower’s utility would decrease. However, we find that this is not always the case. The follower’s utility may either decrease or increase.
Consider the case $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},40,1,(0.01,1),(0.1,1)\rangle$. The case has unique Nash equilibrium, where $\bm{x}_{a}^{*}=(2.884,37.116)$ and $\bm{x}_{b}^{*}=(0.437,0.563)$, resulting in player $a$’s utility being 0.994 and player $b$’s utility being 0.028. In the Stackelberg equilibrium, $\bm{x}_{a}=(1.128,38.872)$ and $\bm{x}_{b}=(0.968,0.032)$, with the leader’s utility rising to 1.005 and the follower’s utility increasing to 0.047. The respective ratios are $\frac{1.005}{0.994}=1.011$ for the leader and $\frac{0.047}{0.028}=1.679$ for the follower.
It is somewhat unexpected that both players’ utilities in the Stackelberg equilibrium exceed their payoffs in the Nash equilibrium. However, this outcome is entirely plausible, as this is a non-zero-sum game. Since the leader values the second battlefield more, they allocate more of their budget to it, minimizing the follower’s share. Meanwhile, the follower values the first battlefield more, and the leader’s strategy incidentally increases the follower’s share of the first battlefield, improving the follower’s overall payoff.
Consider another case $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},0.1,1,(2,1),(10,1)\rangle$. The unique Nash equilibrium is $\bm{x}_{a}^{*}=(0.036,0.064)$ and $\bm{x}_{b}^{*}=(0.738,0.262)$, where player $a$’s utility is 0.289 and player $b$’s utility is 10.338. In the Stackelberg equilibrium, the leader’s allocation is $\bm{x}_{a}=(0.063,0.037)$ and the follower’s is $\bm{x}_{b}=(0.822,0.178)$, yielding a leader utility of 0.315 and a follower utility of 10.118. The respective ratios are $\frac{0.315}{0.289}=1.090$ for the leader and $\frac{10.118}{10.338}=0.979$ for the follower.
In this case, since both players have similar preferences for the battlefields, the leader allocates more budget to the first battlefield to reduce the follower’s share and increase their own payoff. As a result, the follower’s share in the first battlefield decreases, leading to a reduction in their payoff from that battlefield, which in turn causes a decline in their overall payoff.
7 Conclusion and Future Work
We investigate whether, and to what extent, moving first provides an advantage to a player in the Lottery Colonel Blotto game. To address this, we reduce the search space for the follower’s best response strategies and conduct a comparison between the Stackelberg equilibrium and the Nash equilibrium. This comparison is enabled by several insightful constructions and the water-filling understanding of the follower’s best response strategy.
Future research could extend the findings of this study on the Stackelberg version of the Colonel Blotto game in several ways. Additionally, incorporating scenarios with asymmetric information, where players have unequal knowledge about each other’s budgets or battlefield values, can provide insights into real-world applications.
Appendix A OMITTED PROOFS FROM SECTION 3
Lemma 7.
If the leader’s commitment is $\bm{x}_{a}=(x_{a1},\cdots,x_{an})$, the follower’s best response is $\bm{x}_{b}=BR(\bm{x}_{a})=(x_{b1},\cdots,x_{bn})$ in $\mathcal{G}$, and the leader’s commitment is $\bm{x}_{a}^{(1)}=(x_{a1},\cdots,x_{a(r-1)},x_{a(r+1)},\cdots,x_{an},\frac{x_{ar}}{t},\cdots,\frac{x_{ar}}{t})$ in $\mathcal{G}^{(1)}$, we have $\bm{x}_{b}^{(1)}=BR(\bm{x}_{a}^{(1)})=(x_{b1},\cdots,x_{b(r-1)},x_{b(r+1)},\cdots,x_{bn},\frac{x_{br}}{t},\cdots,\frac{x_{br}}{t})$. Furthermore, we have that $u_{a}(\bm{x}_{a}^{(1)},\bm{x}_{b}^{(1)})=u_{a}(\bm{x}_{a},\bm{x}_{b})$.
Proof.
Obviously, there are $n_{1}=n-1+t$ battlefields in $\mathcal{G}^{(1)}$. By Lemma 1, there exists a set $K(\bm{x}_{a})$ in $\mathcal{G}$. Therefore, there are two cases regarding $r$: (1) If $r\notin K(\bm{x}_{a})$, we have $x_{bn+1}^{(1)}=x_{bn+2}^{(1)}=\cdots=x_{bn+t}^{(1)}=0$, and $K(\bm{x}_{a}^{(1)})=K(\bm{x}_{a})$. (2) If $r\in K(\bm{x}_{a})$, we have $x_{bn+1}^{(1)}=x_{bn+2}^{(1)}=\cdots=x_{bn+t}^{(1)}=\frac{x_{br}}{t}$, and $K(\bm{x}_{a}^{(1)})=K(\bm{x}_{a})\backslash\{r\}\cup\{n+1,\cdots,n+t\}$. We substitute $K(\bm{x}_{a}^{(1)})$, $\bm{x}_{b}^{(1)}$ and $\bm{x}_{a}^{(1)}$ into Equation (3), and it is easy to verify that Equation (3) is established.
The aggregate utility from battlefields $\{1,2,\cdots,r-1,r+1,\cdots,n\}$ is equal for players in both $\mathcal{G}$ and $\mathcal{G}^{(1)}$. Consider the last $t$ battlefields in $\mathcal{G}^{(1)}$, we have that
$\sum_{j=n+1}^{n+t}u_{bj}(\bm{x}_{a}^{(1)},\bm{x}_{b}^{(1)})=\sum_{j=n+1}^{n+t}\frac{\frac{x_{br}}{t}}{\frac{x_{ar}}{t}+\frac{x_{br}}{t}}\frac{v_{br}}{t}=u_{br}(\bm{x}_{a},\bm{x}_{b}).$
Similarly, we can also obtain $\sum_{j=n+1}^{n+t}u_{aj}(\bm{x}_{a}^{(1)},\bm{x}_{b}^{(1)})=u_{ar}(\bm{x}_{a},\bm{x}_{b})$. Therefore, $u_{a}(\bm{x}_{a}^{(1)},\bm{x}_{b}^{(1)})=u_{a}(\bm{x}_{a},\bm{x}_{b})$. ∎
Lemma 8.
In $\mathcal{G}^{(2)}$, the leader allocates at least the same amount of resources to the battlefields with higher leader’s valuations in the optimal commitment.
Proof.
We prove it by contradiction. Let $\bm{x}_{a}^{(2)}$ and $\bm{x}_{b}^{(2)}$ denote the leader’s commitment and the follower’s best response respectively. If there are two distinct battlefields $j$ and $j^{\prime}$ satisfying $x_{aj}^{(2)}>x_{aj^{\prime}}^{(2)}$ and $v_{aj}^{(2)}<v_{aj^{\prime}}^{(2)}$. Consider another commitment $\bm{\hat{x}}_{a}^{(2)}$ where $\hat{x}_{aj}^{(2)}=x_{aj^{\prime}}^{(2)}$, $\hat{x}_{aj^{\prime}}^{(2)}=x_{aj}^{(2)}$ and $\hat{x}_{ah}^{(2)}=x_{ah}^{(2)}$ for $\forall h\neq j,j^{\prime}$. By Lemma 1, it is easy to verify that $\bm{\hat{x}}_{b}^{(2)}=BR(\bm{\hat{x}}_{a}^{(2)})$ where $\hat{x}_{bj}^{(2)}=x_{bj^{\prime}}^{(2)}$, $\hat{x}_{bj^{\prime}}^{(2)}=x_{bj}^{(2)}$ and $\hat{x}_{bh}^{(2)}=x_{bh}^{(2)}$ for $\forall h\neq j,j^{\prime}$. In $\hat{\bm{x}}_{a}^{(2)}$, the leader and the follower allocate resources to battlefields other than $j$ and $j^{\prime}$ in the same manner as in $\bm{x}_{a}^{(2)}$; hence, our analysis needs to focus solely on the aggregate utility at battlefields $j$ and $j^{\prime}$. In both $\bm{x}_{a}^{(2)}$ and $\hat{\bm{x}}_{a}^{(2)}$, the difference in the aggregate utility from battlefield $j$ and battlefield $j^{\prime}$ is as follows
$\frac{x_{aj^{\prime}}^{(2)}v_{aj}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}+\frac{x_{aj}^{(2)}v_{aj^{\prime}}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj}^{(2)}v_{aj}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj^{\prime}}^{(2)}v_{aj^{\prime}}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}$ (15)
$=$ $(v_{aj^{\prime}}^{(2)}-v_{aj}^{(2)})\left(\frac{x_{aj}^{(2)}}{x_{aj}^{(2)}+x_{bj}^{(2)}}-\frac{x_{aj^{\prime}}^{(2)}}{x_{aj^{\prime}}^{(2)}+x_{bj^{\prime}}^{(2)}}\right).$
By Equation (3), we obtain $x_{bj}^{(2)}$ and $x_{bj^{\prime}}^{(2)}$ respectively,
$\begin{cases}x_{bj}^{(2)}=\frac{(x_{aj}^{(2)})^{\frac{1}{2}}}{\sum_{l\in K(\bm{x}_{a}^{(2)})}(x_{al}^{(2)})^{\frac{1}{2}}}(x_{b}+\sum_{l\in K(\bm{x}_{a}^{(2)})}x_{al}^{(2)})-x_{aj}^{(2)};\\ x_{bj^{\prime}}^{(2)}=\frac{(x_{aj^{\prime}}^{(2)})^{\frac{1}{2}}}{\sum_{l\in K(\bm{x}_{a}^{(2)})}(x_{al}^{(2)})^{\frac{1}{2}}}(x_{b}+\sum_{l\in K(\bm{x}_{a}^{(2)})}x_{al}^{(2)})-x_{aj^{\prime}}^{(2)}.\\ \end{cases}$
We substitute $x_{bj}^{(2)}$ and $x_{bj^{\prime}}^{(2)}$ into Equation (15), and we find that the value of Equation (15) is positive. Therefore, $u_{a}(\bm{x}_{a}^{(2)},\bm{x}_{b}^{(2)})<u_{a}(\bm{\hat{x}}_{a}^{(2)},\bm{\hat{x}}_{b}^{(2)})$. This immediately implies that for the leader, allocating at least the same amount of resources to the battlefields with higher value within his optimal commitment. ∎
Lemma 9.
In $\mathcal{G}^{(2)}$, the leader allocates equal resources to battlefields with equal leader’s valuations in the optimal commitment.
Proof.
We prove it by contradiction. Let $j$ denote the battlefield with the smallest index such that $x_{aj}^{(2)}<x_{aj+1}^{(2)}$ and $v_{aj}^{(2)}=v_{aj+1}^{(2)}$. Consider another commitment $\bm{\hat{x}}_{a}^{(2)}$ where $\hat{x}_{aj}^{(2)}=\hat{x}_{aj+1}^{(2)}=\frac{1}{2}(x_{aj}^{(2)}+x_{aj+1}^{(2)})$ and $\hat{x}_{ah}^{(2)}=x_{ah}^{(2)}$ for $\forall h\neq j,j+1$. Accordingly, let $\bm{\hat{x}}_{b}^{(2)}=BR(\bm{\hat{x}}_{a}^{(2)})$. Note that if battlefield $j$ is not in $K(\bm{x}_{a}^{(2)})$, then battlefield $j+1$ is also not in $K(\bm{x}_{a}^{(2)})$. Therefore, there are three cases for $K(\bm{x}_{a}^{(2)})$, namely (1) $j\in K(\bm{x}_{a}^{(2)})$ and $j+1\in K(\bm{x}_{a}^{(2)})$, (2) $j\notin K(\bm{x}_{a}^{(2)})$ and $j+1\notin K(\bm{x}_{a}^{(2)})$, and (3) $j\in K(\bm{x}_{a}^{(2)})$ and $j+1\notin K(\bm{x}_{a}^{(2)})$. For each case, we discuss the changes between $K(\bm{x}_{a}^{(2)})$ and $K(\bm{\hat{x}}_{a}^{(2)})$, and compare $u_{a}(\bm{x}_{a}^{(2)},\bm{x}_{b}^{(2)})$ and $u_{a}(\bm{\hat{x}}_{a}^{(2)},\bm{\hat{x}}_{b}^{(2)})$.
For the case (1): according to Lemma 1, we have
$k^{*}=\max\{k\in[n_{2}]:x_{b}+\sum_{l=1}^{j}x_{al}^{(2)}>(x_{aj}^{(2)})^{\frac{1}{2}}\sum_{l=1}^{j}(x_{al}^{(2)})^{\frac{1}{2}},\,\forall j\leq k\}.$ (16)
Due to $j\in K(\bm{x}_{a}^{(2)})$ and $j+1\in K(\bm{x}_{a}^{(2)})$, we immediately obtain the following inequality by Equation (16),
[…]
Due to $x_{aj}^{(2)}<x_{aj+1}^{(2)}$, we can derive the following inequality
[…]
The above inequality implies that $j\in K(\bm{\hat{x}}_{a}^{(2)})$ and $j+1\in K(\bm{\hat{x}}_{a}^{(2)})$.
We analyze the changes in the leader’s utility from $\bm{x}_{a}^{(2)}$ to $\bm{\hat{x}_{a}}^{(2)}$. Consider $\bm{x}_{a}^{(2)\delta}=(x_{a1}^{(2)},\cdots,x_{aj}^{(2)}+\delta,x_{aj+1}^{(2)}-\delta,\cdots,x_{an_{2}}^{(2)})$. If $\delta=\frac{x_{aj+1}^{(2)}-x_{aj}^{(2)}}{2}$, we have $\bm{\hat{x}}_{a}^{(2)}=\bm{x}_{a}^{(2)\delta}$. Let $K(\bm{x}_{a}^{(2)\delta})$ denote the battlefield set corresponding to $\bm{x}_{a}^{(2)\delta}$ and $\bm{x}_{b}^{(2)\delta}=BR(\bm{x}_{a}^{(2)\delta})$. Due to Equation (16), we can know that the set $K(\bm{x}_{a}^{(2)\delta})$ does not expand as $\delta$ increases. Suppose that (i) $\delta_{1}\leq\delta<\delta_{2}$, we have $K(\bm{x}_{a}^{(2)\delta})=K(\bm{x}_{a}^{(2)\delta_{1}})$, and (ii) when $r=|K(\bm{x}_{a}^{(2)\delta_{2}})|$, we have $x_{b}+\sum_{l=1}^{r}x_{al}^{(2)\delta_{2}}=(x_{ar}^{(2)\delta_{2}})^{\frac{1}{2}}\sum_{l=1}^{r}(x_{al}^{(2)\delta_{2}})^{\frac{1}{2}}$, which means that the battlefield $\{r\}=K(\bm{x}_{a}^{(2)\delta})\backslash K(\bm{x}_{a}^{(2)\delta_{2}})$ where $\delta\in[\delta_{1},\delta_{2})$. Note that both $\delta=\delta_{2}$ and $\delta_{1}\leq\delta<\delta_{2}$ work for Equation (16) because of continuity of $\bm{x}_{a}^{(2)}$, which can be viewed as battlefield $r$ is in the set $K(\bm{x}_{a}^{(2)\delta_{2}})$ but $\bm{x}_{br}^{(2)\delta_{2}}$ is exactly zero. According to Equation (2), we have the leader’s aggregate utility $\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}\frac{x_{aj}^{(2)\delta}v_{aj}^{(2)}}{x_{aj}^{(2)\delta}+x_{bj}^{(2)\delta}}$ in set $K(\bm{x}_{a}^{(2)\delta})$, and we substitute the follower’s best response into the leader’s aggregate utility in the set $K(\bm{x}_{a}^{(2)\delta})$ by Equation (3), obtaining that
$\tilde{u}_{a}(\bm{x}_{a}^{(2)\delta})$ $=\left(\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}(\frac{x_{aj}^{(2)\delta}}{v_{bj}^{(2)}})^{\frac{1}{2}}\cdot v_{aj}^{(2)}\right)\left(\frac{\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}(x_{aj}^{(2)\delta}\cdot v_{bj}^{(2)})^{\frac{1}{2}}}{\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}x_{aj}^{(2)\delta}+x_{b}}\right)$
$=\left(\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}(x_{aj}^{(2)\delta})^{\frac{1}{2}}\cdot v_{aj}^{(2)}\right)\cdot\left(\frac{\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}(x_{aj}^{(2)\delta})^{\frac{1}{2}}}{\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}x_{aj}^{(2)\delta}+x_{b}}\right)$
We take the partial derivative of $\tilde{u}_{a}(\bm{x}_{a}^{(2)\delta})$ with regard to $\bm{x}_{aj}^{(2)\delta}$, that is
$\frac{\partial\tilde{u}_{a}(\bm{x}_{a}^{(2)\delta})}{\partial x_{aj}^{(2)\delta}}=$ $\frac{1}{x_{b}+\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}x_{aj}^{(2)\delta}}\bigg{[}\frac{1}{2}\left(\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}(x_{aj}^{(2)\delta})^{\frac{1}{2}}\right)v_{aj}^{(2)}\times(x_{aj}^{(2)\delta})^{-\frac{1}{2}}$
$+(\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}v_{aj}^{(2)}\sqrt{x_{aj}^{(2)\delta}})\times\left(\frac{\frac{1}{2}(x_{aj}^{(2)\delta})^{-\frac{1}{2}}(x_{b}+\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}x_{aj}^{(2)\delta})-\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}(x_{aj}^{(2)\delta})^{\frac{1}{2}}}{x_{b}+\sum_{j\in K(\bm{x}_{a}^{(2)\delta})}x_{aj}^{(2)\delta}}\right)\bigg{]}$
We can find that when $\delta\in[\delta_{1},\delta_{2}]$, $\frac{\partial\tilde{u}_{a}(\bm{x}_{a}^{(2)\delta})}{\partial\bm{x}_{aj}^{(2)\delta}}>\frac{\partial\tilde{u}_{a}(\bm{x}_{a}^{(2)\delta})}{\partial\bm{x}_{aj+1}^{(2)\delta}}$. Therefore, the utility of the leader is improved from $\bm{x}_{a}^{(2)\delta_{1}}$ to $\bm{x}_{a}^{(2)\delta_{2}}$. We keep increasing $\delta$ until it equals $\frac{x_{aj+1}^{(2)}-x_{aj}^{(2)}}{2}$, and the leader’s utility is increasing.
For the case (2): according to Equation (16), we find that the set $K(\bm{x}_{a}^{(2)})=K(\bm{\hat{x}}_{a}^{(2)})$. Given that $x_{aj^{\prime}}^{(2)}=\hat{x}_{aj^{\prime}}^{(2)}$ for $\forall j^{\prime}\in K(\bm{x}_{a}^{(2)})$, it follows that the follower’s best response remains unchanged, consequently, the leader’s utility remains unchanged.
For the case (3): since the set $K(\bm{x}_{a}^{(2)})$ will not expand, the set $K(\bm{x}_{a}^{(2)\delta})$ will undergo following change process: as $\delta$ increases from 0 up to a certain threshold $\delta^{*}$, we have $K(\bm{x}_{a}^{(2)})=K(\bm{x}_{a}^{(2)\delta})$. However, if $\delta$ continues to increase beyond $\delta^{*}$, the set becomes $K(\bm{x}_{a}^{(2)})\backslash\{|K(\bm{x}_{a}^{(2)})|\}=K(\bm{x}_{a}^{(2)\delta})$. During the first stage of changes, the situation is analogous to case (2), and thus the utility remains unchanged. In the second stage of changes, the situation mirror case (1), leading to an increase in utility. This immediately implies that for the leader, allocating equal resources to the battlefields with equal valuation within his optimal commitment. ∎
Lemma 10.
In $\mathcal{G}^{(2)}$, if $v_{ar}^{(2)}=v_{a(r+1)}^{(2)}=\cdots=v_{a(r+t-1)}^{(2)}$ and the leader’s commitment is $\bm{x}_{a}^{(2)}$ where $x_{aj}^{(2)}=x_{ah}^{(2)}$ for $\forall j,h\in\{r,r+1,\cdots,r+t-1\}$. Then in $\mathcal{G}^{(3)}$, for the commitment $\bm{x}_{a}^{(3)}$ where $x_{ar}^{(3)}=tx_{ar}^{(2)}$, $x_{aj}^{(3)}=x_{aj}^{(2)}$ for $\forall j\in\{1,2,\cdots,r-1\}$ and $x_{ah-t+1}^{(3)}=x_{ah}^{(2)}$ for $\forall h\in\{r+t,r+t+1,\cdots,n_{2}\}$, we have $\bm{x}_{b}^{(3)}=BR(\bm{x}_{a}^{(3)})$ where $x_{br}^{(3)}=tx_{br}^{(2)}$ and $x_{bj}^{(3)}=x_{bj}^{(2)}$ for $\forall j\in\{1,2,\cdots,r-1\}$ and $x_{bh-t+1}^{(3)}=x_{bh}^{(2)}$ for $\forall h\in\{r+t,r+t+1,\cdots,n_{2}\}$. Furthermore, $u_{a}(\bm{x}_{a}^{(2)},\bm{x}_{b}^{(2)})=u_{a}(\bm{x}_{a}^{(3)},\bm{x}_{b}^{(3)})$.
Proof.
Similar to the proof method outlined in Lemma 7, there exists a set $K(\bm{x}_{a}^{(2)})$ according to Lemma 1. There are two cases: (1) if $x_{br}^{(2)}=0$, it follows that $x_{br}^{(3)}=0$, and consequently, $K(\bm{x}_{a}^{(2)})=K(\bm{x}_{a}^{(3)})$, and (2) if $x_{br}^{(2)}>0$, it follows that $x_{br}^{(3)}>0$, and consequently, $K(\bm{x}_{a}^{(3)})=K(\bm{x}_{a}^{(2)})\backslash\{r+1,\cdots,r+t-1\}$. Substituting $K(\bm{x}_{a}^{(3)})$, $\bm{x}_{b}^{(3)}$ and $\bm{x}_{a}^{(3)}$ into Equation (3), it is straightforward to confirm that Equation (3) holds. By Equation (1), we find that $\sum_{j=r}^{r+t-1)}u_{aj}(x_{aj}^{(2)},x_{bj}^{(2)})=u_{ar}(x_{ar}^{(3)},x_{br}^{(3)})$, which leads to the conclusion $u_{a}(\bm{x}_{a}^{(2)},\bm{x}_{b}^{(2)})=u_{a}(\bm{x}_{a}^{(3)},\bm{x}_{b}^{(3)})$. ∎
Proof of Theorem 1
Proof.
Let $\mathcal{G}:=\langle\{a,b\},[n],x_{a},x_{b},(v_{aj})_{j=1}^{n},(v_{bj})_{j=1}^{n}\rangle$ denote this game. Then we split every battlefield, and the resulting auxiliary game is denoted as
$\mathcal{G}^{(2)}:=\langle\{a,b\},[n_{2}],x_{a},x_{b},(v_{aj}^{(2)})_{j=1}^{n_{2}},(v_{bj}^{(2)})_{j=1}^{n_{2}}\rangle.$
In $\mathcal{G}^{(2)}$, the follower’s valuation for each battlefield is the same, and for the leader, for a certain battlefield $r$ in $\mathcal{G}$ and the sub-battlefields in $\mathcal{G}^{(2)}$ corresponding to battlefield $r$, the leader’s valuation of these sub-battlefields is the same. Let $\bm{x}^{*}_{a}$ denote the leader’s optimal commitment in $\mathcal{G}$. Let $\bm{x}_{a}^{*(2)}$ represent the leader’s commitment in $\mathcal{G}^{(2)}$ that corresponds to $\bm{x}^{*}_{a}$. Let $\bm{x}_{a}^{(2)*}$ denote the optimal commitment in $\mathcal{G}^{(2)}$. According to Lemma 7, we obtain
$u_{a}(x_{a}^{*},BR(x_{a}^{*}))=u_{a}(\bm{x}_{a}^{*(2)},BR(\bm{x}_{a}^{*(2)}))\leq u_{a}(\bm{x}_{a}^{(2)*},BR(\bm{x}_{a}^{(2)*}))$
Due to $\frac{v_{a1}}{v_{b1}}\leq\frac{v_{a2}}{v_{b2}}\leq\cdots\leq\frac{v_{an}}{v_{bn}}$, it follows that in $\mathcal{G}^{(2)}$, we have $v_{a1}^{(2)}\leq v_{a2}^{(2)}\leq\cdots\leq v_{an_{2}}^{(2)}$. According to Lemma 8, it can be deduced that $x_{a1}^{(2)*}\leq x_{a2}^{(2)*}\leq\cdots\leq x_{an_{2}}^{(2)*}$. According to Lemma 9, we can infer that if $v_{aj}^{(2)}=v_{aj+1}^{(2)}$, then $x_{aj}^{(2)*}=x_{aj+1}^{(2)*}$, for all $j\in[n_{2}-1]$. For any battlefield $r$ in $\mathcal{G}$, we identify the corresponding battlefields in $\mathcal{G}^{(2)}$ and merge these battlefields into a single one. By merging battlefields, we can eventually restore $\mathcal{G}^{(2)}$ to the original game $\mathcal{G}$. Let $\tilde{\bm{x}}_{a}^{*}$ denote the commitment in $\mathcal{G}$ that corresponds to $\bm{x}_{a}^{(2)*}$. Note that $\tilde{\bm{x}}_{a}^{*}$ is a commitment in $\mathcal{G}$. According to Lemma 10, we can deduce that
$u_{a}(\bm{x}_{a}^{(2)*},BR(\bm{x}_{a}^{(2)*}))=u_{a}(\tilde{\bm{x}}_{a}^{*},BR(\tilde{\bm{x}}_{a}^{*}))$
Due to $\bm{x}^{*}_{a}$ is a optimal commitment in $\mathcal{G}$, we have
$u_{a}(\tilde{\bm{x}}_{a}^{*},BR(\tilde{\bm{x}}_{a}^{*}))\leq u_{a}(\bm{x}^{*}_{a},BR(\bm{x}^{*}_{a}))$
Finally, we have following relational expression
$u_{a}(x_{a}^{*},BR(x_{a}^{*}))$ $=u_{a}(\bm{x}_{a}^{*(2)},BR(\bm{x}_{a}^{*(2)}))$
$\leq u_{a}(\bm{x}_{a}^{(2)*},BR(\bm{x}_{a}^{(2)*}))$
$=u_{a}(\tilde{\bm{x}}_{a}^{*},BR(\tilde{\bm{x}}_{a}^{*}))$
$\leq u_{a}(\bm{x}^{*}_{a},BR(\bm{x}^{*}_{a}))$
This implies that
$u_{a}(x_{a}^{*},BR(x_{a}^{*}))$ $=u_{a}(\bm{x}_{a}^{*(2)},BR(\bm{x}_{a}^{*(2)}))$
$=u_{a}(\bm{x}_{a}^{(2)*},BR(\bm{x}_{a}^{(2)*}))$
$=u_{a}(\tilde{\bm{x}}_{a}^{*},BR(\tilde{\bm{x}}_{a}^{*}))$
$=u_{a}(\bm{x}^{*}_{a},BR(\bm{x}^{*}_{a}))$
Thus, the optimal commitment $x_{a}^{*}$ in $\mathcal{G}$, and its corresponding $\bm{x}_{a}^{*(2)}$, is the optimal commitment in $\mathcal{G}^{(2)}$. For battlefield $j$ in $\mathcal{G}$ where $j\in[n]$, the ratio $\frac{x_{aj}^{*}}{v_{bj}}$ remains unchanged before and after the battlefield is split and merged. According to Lemma 8, $x_{a1}^{(2)*}\leq x_{a2}^{(2)*}\leq\cdots\leq x_{an_{2}}^{(2)*}$, we can deduce that for all $j\in[n]$:
$\frac{x_{a1}^{*}}{v_{b1}}\leq\frac{x_{a2}^{*}}{v_{b2}}\leq\cdots\leq\frac{x_{an}^{*}}{v_{bn}}$
Hence,
$\frac{v_{b1}}{x_{a1}^{*}}\geq\frac{v_{b2}}{x_{a2}^{*}}\geq\cdots\geq\frac{v_{bn}}{x_{an}^{*}}$
According to Lemma 1, we can deduce that there are $n$ possible values for $K$, which are $\{1\},\{1,2\},\ldots,$
$\{1,2,3,\ldots,n\}$. In the game $\mathcal{G}$, for $j,h\in[n]$, if $\frac{v_{aj}}{v_{bj}}=\frac{v_{ah}}{v_{bh}}$, then the leader allocates the same amount of resources to the corresponding sub-battlegrounds of battlefields $j$ and $h$, respectively. Therefore, we have $\frac{x_{aj}^{*}}{x_{ah}^{*}}=\frac{v_{aj}}{v_{ah}}$. ∎
Appendix B OMITTED PROOFS FROM SECTION 4
Proof of Lemma 2
Proof.
By Lemma 1, we have $\frac{v_{bj}}{x_{aj}}\leq\frac{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}$ for $j\in\overline{K}$ and rearrange it, we have $x_{aj}\geq\frac{v_{bj}\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}$. Below we show that for a given $K$, $x_{aj}$ cannot be greater than $\frac{v_{bj}\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}$ within the leader’s optimal commitment for $j\in\overline{K}$.
We choose a battlefield $j\in\overline{K}$ such that $j\in\mathop{\arg\max}\limits_{l\in\overline{K}}x_{al}$. Then we transfer an amount $\varepsilon$ of resources from battlefield $j$ to a certain battlefield in $K$, while transferring an amount $\varepsilon$ of resources to other battlefields in $\overline{K}$. The purpose of transferring resources to other battlefields in $\overline{K}$ is to ensure that $K$ remains unchanged when transferring resources to a certain battlefield in $K$. We now prove that after transferring resources, the leader’s utility increases.
If the leader allocates more resources to battlefield $j\in\overline{K}$ than $\frac{v_{bj}\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}$, for this battlefield $j$, we transfer an amount $\varepsilon$ of resources to some battlefield in $K$. If $\varepsilon$ is sufficiently small, $K$ remains unchanged. We demonstrate that the leader’s utility improves after reallocating resources. First, when the leader chooses strategy $\bm{x}_{a}$ and the follower chooses the best response $\bm{x}_{b}$, the leader’s utility on battlefield $j\in K$ is $u_{aj}(\bm{x}_{a},\bm{x}_{b})=\frac{x_{aj}v_{aj}}{x_{aj}+x_{bj}}$. According to Lemma 1, substituting the follower’s best response can yield a utility function solely in terms of $\bm{x}_{a}$,
$u_{aj}(\bm{x}_{a})=\left(\frac{x_{aj}}{v_{bj}}\right)^{\frac{1}{2}}\cdot v_{aj}\cdot\left(\frac{\sum_{h\in K}(x_{ah}\cdot v_{bh})^{\frac{1}{2}}}{x_{b}+\sum_{h\in K}x_{ah}}\right).$
The aggregate utility of the leader over $K$ is
$u_{aK}(\bm{x}_{a})=\sum_{j\in K}u_{aj}(\bm{x}_{a})=\frac{\sum_{h\in K}(x_{ah}v_{bh})^{\frac{1}{2}}}{x_{b}+\sum_{h\in K}x_{ah}}\sum_{j\in K}\left(\frac{x_{aj}}{v_{bj}}\right)^{\frac{1}{2}}v_{aj}.$
If the leader reallocates an infinitesimal amount $\varepsilon$ of resources to battlefield $j\in K$, the partial derivative is given by
$\frac{\partial u_{aK}}{\partial x_{aj}}=\frac{\frac{1}{2}(x_{aj})^{-\frac{1}{2}}(v_{bj})^{\frac{1}{2}}}{\sum_{h\in K}(x_{ah}v_{bh})^{\frac{1}{2}}}-\frac{1}{x_{b}+\sum_{h\in K}x_{ah}}+\frac{\frac{1}{2}\frac{(x_{aj})^{-\frac{1}{2}}}{(v_{bj})^{\frac{1}{2}}}v_{aj}}{\sum_{h\in K}\left(\frac{x_{ah}}{v_{bh}}\right)^{\frac{1}{2}}v_{ah}}.$
Thus, we obtain the following expression
$x_{aj}\frac{\partial u_{aK}}{\partial x_{aj}}=\frac{\frac{1}{2}(x_{aj})^{\frac{1}{2}}(v_{bj})^{\frac{1}{2}}}{\sum_{h\in K}(x_{ah}v_{bh})^{\frac{1}{2}}}-\frac{x_{aj}}{x_{b}+\sum_{h\in K}x_{ah}}+\frac{\frac{1}{2}\frac{(x_{aj})^{\frac{1}{2}}}{(v_{bj})^{\frac{1}{2}}}v_{aj}}{\sum_{h\in K}\left(\frac{x_{ah}}{v_{bh}}\right)^{\frac{1}{2}}v_{ah}}.$
Further, we derive
$\sum_{j\in K}x_{aj}\frac{\partial u_{aK}}{\partial x_{aj}}=1-\frac{\sum_{h\in K}x_{ah}}{x_{b}+\sum_{h\in K}x_{ah}}>0.$
This implies that within the first $K$ battlefields, there must exist a battlefield $r\in K$ such that $\frac{\partial u_{aK}}{\partial x_{ar}}>0$. In other words, if there exists a battlefield $j\in\overline{K}$ where the leader allocates more resources than $\frac{v_{bj}\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}$, the leader can transfer $\varepsilon$ resources from this battlefield to battlefield $r$, thereby strictly increasing the leader’s utility. Consequently, for all $j\in\overline{K}$, the resources allocated by the leader are precisely equal to $\frac{v_{bj}\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}$. ∎
Proof of Lemma 3
Proof.
Let $|K(\bm{x}_{a})|=k$, $\Delta\bm{x}=(\Delta x_{j})_{j=1}^{k}$. Consider another commitment $\bm{x^{\prime}}_{a}=((x_{aj}-\Delta x_{j})_{j=1}^{k},x_{a(k+1)},\cdots,x_{an})$ and $\bm{x^{\prime}}_{b}=BR(\bm{x^{\prime}}_{a})$, and the corresponding battlefield set is $K(\bm{x^{\prime}}_{a})$. Note that $\bm{x}_{a}$ is the optimal commitment, and let $\bm{x}_{b}=BR(\bm{x}_{a})$, therefore, we have that $u_{a}(\bm{x^{\prime}}_{a},\bm{x^{\prime}}_{b})\leq u_{a}(\bm{x}_{a},\bm{x}_{b})$. By Lemma 1, we can know that the follower’s best response is a continuous function of $\bm{x}_{a}$. Note that $x_{bj}>0$ for $j\in K(\bm{x}_{a})$. Therefore, as long as $\Delta\bm{x}$ is small enough, we still have that $x_{bj}>0$, $\forall j\in K(\bm{x}_{a})$. There are two cases: (1) $K(\bm{x}_{a})=K(\bm{x^{\prime}}_{a})$, and (2) $K(\bm{x}_{a})\subsetneq K(\bm{x^{\prime}}_{a})$.
If $K(\bm{x^{\prime}}_{a})$ is greater than $K(\bm{x}_{a})$, the follower must choose some battlefields in $\overline{K(\bm{x}_{a})}$. Note that the action profile $(\bm{x}_{a},\bm{x}_{b})$, the battlefields in the set $\overline{K(\bm{x}_{a})}$ have the same marginal utility for the follower, then the follower will simultaneously invest resources in these battlefields. Consequently, $K(\bm{x^{\prime}}_{a})=[n]$ when $K(\bm{x^{\prime}}_{a})$ is greater than $K(\bm{x}_{a})$. Now we only discuss the case (1) and the case (2) is similar to the case (1), thus the case (2) can be omitted.
First, we note that $\bm{x}_{a}$ is the optimal commitment. When the leader makes the commitment $\bm{x^{\prime}}_{a}$, the leader’s utility do not increase. Therefore, we have the following inequality
$\left(\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{a1}},\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{a2}},\cdots,\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{ak}}\right)\cdot\Delta\bm{x}\leq 0.$
Due to the set $K(\bm{x}_{a})$ does not changed and the battlefields in the set $\overline{K(\bm{x}_{a})}$ does not enter the set $K(\bm{x}_{a})$, according to Lemma 1, we have that $\sum_{h\in K(\bm{x}_{a})}(x_{ah}^{\prime}v_{bh})^{\frac{1}{2}}\leq\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}$. Let $L(\bm{x}_{a})=\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}$. For any $\Delta\bm{x}$ satisfying $(1,1,\cdots,1)\cdot\Delta\bm{x}=0$, which means that the amount of resources allocated by the leader remains unchanged in the set $K(\bm{x}_{a})$, therefore, we have inequality $\left(\frac{\partial L(\bm{x}_{a})}{\partial x_{a1}},\frac{\partial L(\bm{x}_{a})}{\partial x_{a2}},\cdots,\frac{\partial L(\bm{x}_{a})}{\partial x_{ak}}\right)\cdot\Delta\bm{x}\leq 0$ that can be induced by the fact that $\sum_{h\in K(\bm{x}_{a})}(x_{ah}^{\prime}v_{bh})^{\frac{1}{2}}\leq\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}$. Furthermore we have that $\left(\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{aj}}\right)_{j\in K(\bm{x}_{a})}$ is a linear combination of $\left(\frac{\partial L(\bm{x}_{a})}{\partial x_{aj}}\right)_{j\in K(\bm{x}_{a})}$ and $(1,1,\cdots,1)$. We take the partial derivative $\hat{u}_{a}(\bm{x}_{a})$ and $L(\bm{x}_{a})$ with regard to $x_{aj}$ for $j\in K(\bm{x}_{a})$ respectively,
$\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{aj}}=$ $\frac{1}{2}\frac{v_{aj}}{\sqrt{v_{bj}x_{aj}}}\frac{\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}}{x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah}}+\left(\sum_{h\in K(\bm{x}_{a})}\frac{v_{ah}\sqrt{x_{ah}}}{\sqrt{v_{bh}}}\right)\frac{\frac{1}{2}\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}(x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah})-\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}}{(x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah})^{2}}$
$=$ $\left(\frac{1}{2}\frac{\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}}{x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah}}\frac{v_{aj}}{v_{bj}}+\frac{1}{2}\frac{\sum_{h\in K(\bm{x}_{a})}\frac{v_{ah}\sqrt{x_{ah}}}{\sqrt{v_{bh}}}}{x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah}}\right)\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}-\frac{\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}\sum_{h\in K(\bm{x}_{a})}\frac{v_{ah}\sqrt{x_{ah}}}{\sqrt{v_{bh}}}}{(x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah})^{2}},$
$\frac{\partial L(\bm{x}_{a})}{\partial x_{aj}}=\frac{\partial\sum_{h\in K(\bm{x})}(x_{ah}v_{bh})^{\frac{1}{2}}}{\partial x_{aj}}=\frac{1}{2}\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}.$
For convenience, we define
$\begin{cases}C_{1}=\frac{1}{2}\frac{\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}}{x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah}};\\ C_{2}=\frac{1}{2}\frac{\sum_{h\in K(\bm{x}_{a})}\frac{v_{ah}\sqrt{x_{ah}}}{\sqrt{v_{bh}}}}{x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah}};\\ C_{3}=\frac{\sum_{h\in K(\bm{x}_{a})}(x_{ah}v_{bh})^{\frac{1}{2}}\sum_{h\in K(\bm{x}_{a})}\frac{v_{ah}\sqrt{x_{ah}}}{\sqrt{v_{bh}}}}{(x_{b}+\sum_{h\in K(\bm{x}_{a})}x_{ah})^{2}},\end{cases}$
we have
$\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{aj}}=(C_{1}\frac{v_{aj}}{v_{bj}}+C_{2})\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}-C_{3}.$ (17)
$\left(\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{aj}}\right)_{j\in K(\bm{x}_{a})}$ is a linear combination of $\left(\frac{\partial L(\bm{x}_{a})}{\partial x_{aj}}\right)_{j\in K(\bm{x}_{a})}$ and $(1,1,\cdots,1)$, we have
$\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{aj}}=C_{4}\frac{\partial L(\bm{x}_{a})}{\partial x_{aj}}+C_{5}.$ (18)
We substitute $\frac{\partial L(\bm{x}_{a})}{\partial x_{aj}}=\frac{1}{2}\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}$ into Equation (18) and obtain
$\frac{\partial\hat{u}_{a}(\bm{x}_{a})}{\partial x_{aj}}=\frac{1}{2}C_{4}\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}+C_{5}$
Therefore, we have
$(C_{1}\frac{v_{aj}}{v_{bj}}+C_{2})\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}-C_{3}=C_{4}\frac{\partial L(\bm{x}_{a})}{\partial x_{aj}}+C_{5}.$ (19)
Arranging Equation (19) can yield
$\frac{v_{aj}}{v_{bj}}\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}=\frac{C_{4}-C_{2}}{C_{1}}\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}+\frac{C_{5}+C_{3}}{C_{1}}$ (20)
Let $\alpha=\frac{C_{4}-C_{2}}{C_{1}}$ and $\beta=\frac{C_{5}+C_{3}}{C_{1}}$, we have $\frac{v_{aj}}{v_{bj}}\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}=\alpha\frac{\sqrt{v_{bj}}}{\sqrt{x_{aj}}}+\beta$. Rearranging it, we can obtain $\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}}=\sqrt{x_{aj}}\beta$ for $\forall j\in K$. ∎
The rest of the Proof of Theorem 2
Proof.
Let us analyze the first subcase of Case 2, i.e., Case 2.1. Because of $K=[n]$, we have $\frac{1}{\beta^{2}}\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}=x_{a}$. It follows that $\beta^{2}=\frac{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{x_{a}}$. According to Problem (11), we have
$\max_{\bm{x}_{a}\in\bm{X}_{a}}\quad$ $\hat{u}_{a}(\bm{x}_{a})=\frac{\left(\sum_{j\in[n]}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in[n]}(v_{aj}-\alpha v_{bj})\right)}{\frac{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{x_{a}}x_{b}+\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$ (21)
s.t. $\frac{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{x_{a}}>0.$
We need to find a suitable parameter $\alpha$ that maximizes the leader’s utility $\hat{u}_{a}(\bm{x}_{a})$. Therefore, the optimal commitment is as follows
$x_{aj}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\cdot\frac{x_{a}}{\sum_{j\in[n]}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}},\;\forall j\in[n].$ (22)
To find a suitable parameter $\alpha$, we rewrite the Problem 21 as follows
$\max_{\bm{x}_{a}\in\bm{X}_{a}}\hat{u}_{a}(\bm{x}_{a})=\frac{\left((\sum_{j\in[n]}v_{aj})^{2}-(\sum_{j\in[n]}v_{bj})(\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}})\right)\alpha}{\alpha^{2}(\sum_{j\in[n]}v_{bj})-2\alpha(\sum_{j\in[n]}v_{aj})+\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}}}$ (23)
Note that $(\sum_{j\in[n]}v_{aj})^{2}-(\sum_{j\in[n]}v_{bj})(\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}})<0$, we choose the $\alpha$ with negative value to maximize the leader’s utility. Solve the Problem 23, we have $\alpha=-\sqrt{\frac{1}{\sum_{j\in[n]}v_{bj}}\sum_{j\in[n]}\frac{v_{aj}^{2}}{v_{bj}}}$. Substitute $\alpha$ into Equation (22), the closed-form solution is obtained, that is
$x_{aj}=\frac{\left(\frac{v_{aj}}{\sqrt{v_{bj}}}+\left(\frac{\sum_{h=1}^{n}\frac{(v_{ah})^{2}}{v_{bh}}}{\sum_{h=1}^{n}v_{bh}}\right)^{\frac{1}{2}}\sqrt{v_{bj}}\right)^{2}}{\sum_{l=1}^{n}\left(\frac{v_{al}}{\sqrt{v_{bl}}}+\left(\frac{\sum_{h=1}^{n}\frac{(v_{ah})^{2}}{v_{bh}}}{\sum_{h=1}^{n}v_{bh}}\right)^{\frac{1}{2}}\sqrt{v_{bl}}\right)^{2}}\cdot x_{a}.$
Finally, let us analyse the sub-case 2.1. Because of $K\subsetneq[n]$, we can derive following equation
$x_{a}=\frac{1}{\beta^{2}}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}+(\sum_{j\in\overline{K}}v_{bj})\frac{\left[\beta x_{b}+\frac{1}{\beta}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\right]^{2}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}.$ (24)
Let $y=\beta^{2}$, we have $x_{aj}^{\frac{1}{2}}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\beta}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\sqrt{y}}$. Substituting $x_{aj}^{\frac{1}{2}}=(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\sqrt{y}}$ into the objective function of Problem 11, we obtain the leader’s utility, that is
$\hat{u}_{a}(\bm{x}_{a})$ $=\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)\frac{1}{\sqrt{y}}v_{aj}\right)\frac{\sum_{j\in K}(v_{aj}-\alpha v_{bj})\frac{1}{\sqrt{y}}}{\frac{1}{y}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})^{2}+x_{b}}$
$=\frac{1}{y}\cdot\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{x_{b}+\frac{1}{y}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$
$=\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{yx_{b}+\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$ (25)
Obviously, $\hat{u}_{a}(\bm{x}_{a})$ increases as $y$ decreases for a given $\alpha$. We can rewrite Equation (24) as
$x_{a}=y^{-1}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}+(\sum_{j\in\overline{K}}v_{bj})\frac{y\left(x_{b}+y^{-1}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\right)^{2}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}.$ (26)
For convenience, we define
$\begin{cases}A_{1}=\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\\ A_{2}=\frac{\sum_{j\in\overline{K}}v_{bj}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}\\ \end{cases}$
Therefore, we have
$y^{-1}A_{1}+A_{2}y\left(x_{b}+y^{-1}A_{1}\right)^{2}=x_{a}.$
Furthermore, it can be rearranged as
$A_{2}x_{b}^{2}y^{2}+(2A_{1}A_{2}x_{b}-x_{a})y+A_{1}+A_{2}A_{1}^{2}=0,$
which is a quadratic equation about $y$. Let $y_{1}$ and $y_{2}$ be roots of this equation, respectively. We have $y_{1}y_{2}>0$, which implies $y_{1}>0$ and $y_{2}>0$. For a given $\alpha$, we choose a smaller root $\hat{y}=\min\{y_{1},y_{2}\}$ to maximize the leader’s utility. Thus, we have
$\hat{y}$ $=\frac{(x_{a}-2A_{1}A_{2}x_{b})-\sqrt{(2A_{1}A_{2}x_{b}-x_{a})^{2}-4A_{2}x_{b}^{2}(A_{1}+A_{2}A_{1}^{2})}}{2A_{2}x_{b}^{2}}$
$=\frac{(x_{a}-2A_{1}A_{2}x_{b})-\sqrt{x_{a}^{2}-4A_{1}A_{2}x_{a}x_{b}-4A_{1}A_{2}x_{b}^{2}}}{2A_{2}x_{b}^{2}}$
$=\frac{\left(x_{a}-2\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\right)(\sum_{j\in\overline{K}}v_{bj})}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}x_{b}\right)}{2\frac{\sum_{j\in\overline{K}}v_{bj}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}x_{b}^{2}}-\frac{\sqrt{x_{a}^{2}-4\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\right)(\sum_{j\in\overline{K}}v_{bj})}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}x_{b}(x_{a}+x_{b})}}{2\frac{\sum_{j\in\overline{K}}v_{bj}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}x_{b}^{2}}$
After simplification, we obtain
$\hat{y}=\frac{B_{1}\alpha^{2}+B_{2}\alpha+B_{3}-(v_{aK}-v_{bK}\alpha)\sqrt{B_{4}\alpha^{2}+B_{5}\alpha+B_{6}}}{2x_{b}^{2}v_{b\overline{K}}}$
When $B_{4}\alpha^{2}+B_{5}\alpha+B_{6}\geq 0$, $y$ has real number solutions. Substitute $\hat{y}$ into Equation (25), and obtain the following optimization problem
$\max_{\bm{x_{a}}\in\bm{X}_{a}}\quad$ $\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)v_{aj}\right)\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{\hat{y}x_{b}+\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$ (27)
s.t. $x_{a}^{2}-4\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\right)(\sum_{j\in\overline{K}}v_{bj})}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}x_{b}(x_{a}+x_{b})\geq 0.$
We need to find a suitable parameter $\alpha$ that maximizes the leader’s utility. Therefore, the optimal commitment is as follows
$x_{aj}=\begin{cases}\frac{2x_{b}^{2}v_{b\overline{K}}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{B_{1}\alpha^{2}+B_{2}\alpha+B_{3}-(v_{aK}-v_{bK}\alpha)\sqrt{B_{4}\alpha^{2}+B_{5}\alpha+B_{6}}},\;j\in K;\\ \frac{v_{bj}\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},\;j\in\overline{K}.\end{cases}$
Proof of Lemma 4
Proof.
We prove it by contradiction. Suppose the leader does not allocate resources in some battlefields. Let $Q=\{1,2,\cdots,q\}$ be a battlefield set where the leader allocates positive resources, $q<n$. Let $\overline{Q}=\{q+1,\cdots,n\}$ be a battlefield set where the leader does not allocate resources. Without loss of generality, we assume that $\frac{v_{b1}}{x_{a1}}\geq\cdots\geq\frac{v_{bq}}{x_{aq}}$. Let $K$ denote the battlefield set where the follower allocates positive resources in his best response, $|K|=k$. Then we have $\frac{v_{bk}}{x_{ak}}>\frac{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{(x_{b}+\sum_{l\in K}x_{al})^{2}}$ according to Lemma 1.
If $k+1\leq q$, we have
$\frac{v_{b(k+1)}}{x_{a(k+1)}}\leq\frac{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}+(x_{a(k+1)}^{\frac{1}{2}}v_{b(k+1)}^{\frac{1}{2}})\right)^{2}}{(x_{b}+\sum_{l\in K}x_{al}+x_{a(k+1)})^{2}}.$
We define the another commitment $\hat{\bm{x}}_{a}(\varepsilon)$ such that $\hat{x}_{al}(\varepsilon)=(1-\varepsilon)x_{al}$ for $1\leq l\leq k$, $\hat{x}_{an}(\varepsilon)=\varepsilon\sum_{l\in K}x_{al}$, and $\hat{x}_{al}(\varepsilon)=x_{al}$ for $k+1\leq l\leq n-1$. When $\varepsilon$ is a small enough positive number, according to Lemma 1, we can easily verify that the following inequality holds
$\frac{v_{bk}}{\hat{x}_{ak}(\varepsilon)}>\frac{\left(\sum_{l\in K}(\hat{x}_{al}(\varepsilon)v_{bl})^{\frac{1}{2}}+(\hat{x}_{an}(\varepsilon)v_{bn})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l\in K}\hat{x}_{al}(\varepsilon)+\hat{x}_{an}(\varepsilon)\right)^{2}}.$
It implies that $K(\hat{\bm{x}}_{a}(\varepsilon))=\{n\}\cup K(\bm{x}_{a})$. Thus, we have $\frac{v_{bn}}{\hat{x}_{an}(\varepsilon)}\geq\frac{v_{b1}}{\hat{x}_{a1}(\varepsilon)}\geq\cdots\geq\frac{v_{bk}}{\hat{x}_{ak}(\varepsilon)}$. On the other hand, because of $k+1\leq q$, we can easily verify that
$\frac{v_{b(k+1)}}{\hat{x}_{a(k+1)}(\varepsilon)}\leq\frac{\left(\sum_{l\in K}(\hat{x}_{al}(\varepsilon)v_{bl})^{\frac{1}{2}}+(\hat{x}_{an}(\varepsilon)v_{bn})^{\frac{1}{2}}+(\hat{x}_{a(k+1)}(\varepsilon)v_{b(k+1)})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l\in K}\hat{x}_{al}(\varepsilon)+\hat{x}_{an}(\varepsilon)\right)^{2}}.$
Finally, we prove that the leader’s utility increases when the leader makes commitment $\hat{\bm{x}}_{a}(\varepsilon)$ compared to $\bm{x}_{a}$.
$u_{a}(\hat{\bm{x}}_{a}(\varepsilon))=$ $\left(\sum_{j\in K}(\frac{\hat{x}_{aj}(\varepsilon)}{v_{bj}})^{\frac{1}{2}}v_{aj}+(\frac{\hat{x}_{an}(\varepsilon)}{v_{bn}})^{\frac{1}{2}}v_{an}\right)\times\left(\frac{\sum_{l\in K}(\hat{x}_{al}(\varepsilon)v_{bl})^{\frac{1}{2}}+(\hat{x}_{an}(\varepsilon)v_{bn})^{\frac{1}{2}}}{\sum_{l\in K}\hat{x}_{al}(\varepsilon)+\hat{x}_{an}(\varepsilon)+x_{b}}\right)$
$=$ $\left((1-\varepsilon)^{\frac{1}{2}}\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}v_{aj}+\frac{(\varepsilon x_{aK})^{\frac{1}{2}}}{v_{bn}^{\frac{1}{2}}}v_{an}\right)\times\left(\frac{(1-\varepsilon)^{\frac{1}{2}}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}+\varepsilon^{\frac{1}{2}}x_{aK}^{\frac{1}{2}}v_{bn}^{\frac{1}{2}}}{x_{K}+x_{b}}\right)$
$=$ $\frac{1}{x_{aK}+x_{b}}\Bigg{[}(1-\varepsilon)\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}v_{aj}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}+\varepsilon x_{aK}v_{an}+(\varepsilon-\varepsilon^{2})^{\frac{1}{2}}\Big{[}(x_{aK}v_{bn})^{\frac{1}{2}}\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}v_{aj}\Big{]}+$
$(\varepsilon-\varepsilon^{2})^{\frac{1}{2}}(\frac{x_{aK}}{v_{bn}})^{\frac{1}{2}}v_{an}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}\Bigg{]}$
where $x_{aK}=\sum_{j\in K}x_{aj}$. The change of the leader’s utility is that
$u_{a}(\hat{\bm{x}}_{a}(\varepsilon))-u_{a}(\bm{x}_{a})=$ $\frac{\varepsilon}{x_{aK}+x_{b}}\Bigg{[}-\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}+x_{aK}v_{an}+(\frac{1-\varepsilon}{\varepsilon})^{\frac{1}{2}}\Big{[}(x_{aK}v_{bn})^{\frac{1}{2}}\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}+$
$(\frac{x_{aK}}{v_{bn}})^{\frac{1}{2}}v_{an}\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}\Big{]}\Bigg{]}$
When $\varepsilon$ is small enough, we have $u_{a}(\hat{\bm{x}}_{a}(\varepsilon))-u_{a}(\bm{x}_{a})>0$. If there is a battlefield where the leader does not allocate resources, then the leader can transfer $\varepsilon$ resources to that battlefield, thereby increasing the utility of the leader. The leader allocates positive resources across all battlefields. ∎
Appendix C OMITTED PROOFS FROM SECTION 5
Proof of Lemma 5
Proof.
Previous works Kim et al. [2018]; Li and Zheng [2022] have established that if the strategy profile $(\bm{x}_{a},\bm{x}_{b})$ is a NE, then $x_{ij}>0$, for $\forall i\in\{a,b\}$, $\forall j\in[n]$. Hence, if $(\bm{x}_{a},BR(\bm{x}_{a}))$ constitutes a NE, we must have $K(\bm{x}_{a})=[n]$. Therefore, we examine the situation where the leader’s commitment results in the follower having positive best responses across all battlefields.
According to Equation (3), if $\hat{\bm{x}}_{b}$ is the best response to $\hat{\bm{x}}_{a}$ and $\hat{\bm{x}}_{a}$ is also the best response to $\hat{\bm{x}}_{b}$, then we have
$\begin{cases}\hat{\bm{x}}_{b}=\left(\frac{(\hat{x}_{aj}v_{bj})^{\frac{1}{2}}(x_{b}+x_{a})}{\sum_{j^{\prime}=1}^{n}(\hat{x}_{aj^{\prime}}v_{bj^{\prime}})^{\frac{1}{2}}}-\hat{x}_{aj}\right)_{j=1,2,\cdots,n};\\ \hat{\bm{x}}_{a}=\left(\frac{(\hat{x}_{bj}v_{aj})^{\frac{1}{2}}(x_{a}+x_{b})}{\sum_{k=1}^{n}(\hat{x}_{bk}v_{ak})^{\frac{1}{2}}}-\hat{x}_{bj}\right)_{j=1,2,\cdots,n}.\end{cases}$ (28)
By Equation (28), we have the following equation
$\frac{(\hat{x}_{aj}v_{bj})^{\frac{1}{2}}(x_{a}+x_{b})}{\sum_{k=1}^{n}(\hat{x}_{ak}\cdot v_{bk})^{\frac{1}{2}}}=\hat{x}_{aj}+\hat{x}_{bj}=\frac{(\hat{x}_{bj}v_{aj})^{\frac{1}{2}}(x_{a}+x_{b})}{\sum_{k=1}^{n}(\hat{x}_{bk}\cdot v_{ak})^{\frac{1}{2}}},\forall j\in[n].$ (29)
In Equation (29), the first equality holds by using Lemma 1 and the fact that $\hat{\bm{x}}_{b}$ is best response to $\hat{\bm{x}}_{a}$. The second equality holds by using Lemma 1 and the fact that $\hat{\bm{x}}_{a}$ is also best response to $\hat{\bm{x}}_{b}$.
By Equation (29), we have $\frac{(\hat{x}_{aj}\cdot v_{bj})^{\frac{1}{2}}}{(\hat{x}_{bj}\cdot v_{aj})^{\frac{1}{2}}}=\frac{\sum_{k=1}^{n}(\hat{x}_{ak}\cdot v_{bk})^{\frac{1}{2}}}{\sum_{k=1}^{n}(\hat{x}_{bk}\cdot v_{ak})^{\frac{1}{2}}}$, implying $\forall j\in[n]$, $\frac{\hat{x}_{aj}\cdot v_{bj}}{\hat{x}_{bj}\cdot v_{aj}}$ is a constant. We substitute $\hat{x}_{bj}$ into Equation (29) yielding
$\frac{\hat{x}_{aj}\cdot v_{bj}}{\hat{x}_{bj}\cdot v_{aj}}=\frac{\hat{x}_{aj}\cdot v_{bj}}{\left[\frac{(\hat{x}_{aj}\cdot v_{bj})^{\frac{1}{2}}\cdot(x_{a}+x_{b})}{\sum_{k=1}^{n}(\hat{x}_{ak}\cdot v_{bk})^{\frac{1}{2}}}-\hat{x}_{aj}\right]\cdot v_{aj}}=\frac{1}{\frac{x_{a}+x_{b}}{\sum_{k=1}^{n}(\hat{x}_{ak}v_{bk})^{\frac{1}{2}}}\cdot\frac{v_{aj}}{(\hat{x}_{aj}v_{bj})^{\frac{1}{2}}}-\frac{v_{aj}}{v_{bj}}}.$ (30)
According to Lemma 3, we have $\sqrt{\hat{x}_{aj}}=\frac{1}{\beta}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\sqrt{v_{bj}}\alpha)$. Let $\gamma=\frac{1}{\beta}$ and $r=-\alpha$, we have $\sqrt{\hat{x}_{aj}}=\gamma(\frac{v_{aj}}{\sqrt{v_{bj}}}+r\sqrt{v_{bj}})$. Substituting it into Equation (30) yields
$\frac{\hat{x}_{bj}\cdot v_{aj}}{\hat{x}_{aj}\cdot v_{bj}}$ $=\frac{x_{a}+x_{b}}{\sum_{k=1}^{n}(\hat{x}_{ak}\cdot v_{bk})^{\frac{1}{2}}}\cdot\frac{v_{aj}}{\gamma(\frac{v_{aj}}{\sqrt{v_{bj}}}+r\sqrt{v_{bj}})(v_{bj})^{\frac{1}{2}}}-\frac{v_{aj}}{v_{bj}}$ (31)
$=\frac{x_{a}+x_{b}}{\sum_{k=1}^{n}(\hat{x}_{ak}\cdot v_{bk})^{\frac{1}{2}}}\cdot\frac{v_{aj}}{\gamma(v_{aj}+rv_{bj})}-\frac{v_{aj}}{v_{bj}}$
$=\frac{x_{a}+x_{b}}{\sum_{k=1}^{n}(\hat{x}_{ak}\cdot v_{bk})^{\frac{1}{2}}}\cdot\frac{1}{\gamma(1+r\frac{v_{bj}}{v_{aj}})}-\frac{v_{aj}}{v_{bj}}.$
Let $t=\frac{v_{aj}}{v_{bj}}$, $D=\frac{x_{a}+x_{b}}{\sum_{k=1}^{n}(\hat{x}_{ak}\cdot v_{bk})^{\frac{1}{2}}}$, and let $E=\frac{\hat{x}_{bj}\cdot v_{aj}}{\hat{x}_{aj}\cdot v_{bj}}$, then Equation (31) can be rewritten as
$E=D\cdot\frac{1}{\gamma(1+\frac{r}{t})}-t.$ (32)
We rearrange Equation (32) and obtain $\gamma t^{2}+(\gamma E+\gamma r-D)t+\gamma Er=0$. The terms $r,D,E,\gamma$ are constants independent of $t$, therefore we can view $t$ as a variable. Equation (32) is a quadratic equation with regard to $t$. Therefore, for any profile $(D,\gamma,r,E)$, there are at most two solutions for $t$ in the Equation (32), which means that $t=\frac{v_{aj}}{v_{bj}}$ may have two different values at most such that the optimal commitment and its best response constitute a Nash equilibrium. ∎
Proof of Theorem 3
Proof.
Let $\bm{x}_{a}$ denote the leader’s optimal commitment and $\bm{x}_{b}$ denote the follower’s best response. For the first case, it is easy to verify that $\forall j\in[n]$, $x_{aj}=\frac{v_{aj}x_{a}}{\sum_{h\in[n]}v_{ah}}$ and $x_{bj}=\frac{v_{bj}x_{b}}{\sum_{h\in[n]}v_{bh}}$ can establish a NE.
For the second case, we consider the new game after merging $n$ battlefields into two battlefields. Let $\hat{\bm{x}}_{a}=(\hat{x}_{a1},\hat{x}_{a2})$ and $\hat{\bm{x}}_{b}=(\hat{x}_{b1},\hat{x}_{b2})$ denote the leader’s optimal commitment strategy and the follower’s best response respectively. By Theorem 2, we can get the leader’s optimal commitment, that is
$\begin{cases}\hat{x}_{a1}=\frac{t_{1}}{t_{1}+t_{2}}x_{a};\\ \hat{x}_{a2}=\frac{t_{2}}{t_{1}+t_{2}}x_{a}.\end{cases}$
If $\hat{\bm{x}}_{a}=(\hat{x}_{a1},\hat{x}_{a2})$ is also a leader’s Nash equilibrium strategy, we have
$\frac{\hat{x}_{a1}v_{bM}}{\hat{x}_{b1}v_{aM}}=\frac{\hat{x}_{a2}v_{b\overline{M}}}{\hat{x}_{b2}v_{a\overline{M}}}.$ (33)
By Lemma 1, we have $\hat{x}_{bj}=\frac{\left(\hat{x}_{aj}v_{bM}\right)^{\frac{1}{2}}(x_{b}+x_{a})}{\left(\hat{x}_{a1}v_{bM}\right)^{\frac{1}{2}}+\left(\hat{x}_{a2}v_{b\overline{M}}\right)^{\frac{1}{2}}}-\hat{x}_{aj}$, and we substitute it into Equation (33). After some calculations we get
$\frac{x_{a}}{x_{b}}=\frac{t_{4}}{t_{3}}.$ (34)
Therefore, we can know $\hat{x}_{a1}>0$, $\hat{x}_{a2}>0$, and Equation (33) holds when Equation (34) is established, and $\hat{x}_{b1}>0$, $\hat{x}_{b2}>0$.
Finally, we need to verify that when Equation (34) is established, $\hat{\bm{x}}_{a}=(\hat{x}_{a1},\hat{x}_{a2})$ we solved is indeed optimal. By Lemma 1, we know that there are two cases in the follower’s best response: (1) positive resources are invested in all battlefields in the follower’s best response, and (2) there is one battlefield with no resources invested, and another battlefield with all resources invested in the follower’s best response. Without loss of generality, let $(\bm{x^{\prime}}_{a1},\bm{x^{\prime}}_{a2})$ be the another commitment such that $\bm{x^{\prime}}_{b1}=x_{b}$ and $\bm{x^{\prime}}_{b2}=0$. Because of $\bm{x^{\prime}}_{b2}=0$, the leader must invest more resources in the battlefield 2. Thus, we have $x^{\prime}_{a2}>\hat{x}_{a2}$ and $x^{\prime}_{a1}<\hat{x}_{a1}$. By Equation (11), we can calculate the partial derivative of $u_{a}$ with respect to $\bm{x}_{a}$, that is
$\frac{\partial u_{a}}{\partial x_{a1}}=\frac{1}{2(x_{a}+x_{b})}\left(2v_{aM}+\left(\frac{v_{aM}v_{b\overline{M}}^{\frac{1}{2}}}{v_{bM}^{\frac{1}{2}}}+\frac{v_{a\overline{M}}v_{bM}^{\frac{1}{2}}}{v_{b\overline{M}}^{\frac{1}{2}}}\right)\frac{x_{a2}^{\frac{1}{2}}}{x_{a1}^{\frac{1}{2}}}\right).$ (35)
Then we have $\frac{\partial u_{a}}{\partial x^{\prime}_{a1}}>\frac{\partial u_{a}}{\partial\hat{x}_{a1}}$ and $\frac{\partial u_{a}}{\partial x^{\prime}_{a2}}<\frac{\partial u_{a}}{\partial\hat{x}_{a2}}$. Due to $\frac{\partial u_{a}}{\partial\hat{x}_{a1}}=\frac{\partial u_{a}}{\partial\hat{x}_{a2}}$, then $\frac{\partial u_{a}}{\partial x^{\prime}_{a1}}>\frac{\partial u_{a}}{\partial x^{\prime}_{a2}}$. Therefore, the leader is willing to increase $x^{\prime}_{a1}$ and decrease $x^{\prime}_{a2}$. We can conclude that $\bm{x^{\prime}}_{a}$ is not a optimal commitment and $\bm{\hat{x}}_{a}$ is indeed an optimal commitment.
It is worth noting that when $\frac{v_{aj}}{v_{bj}}=c$ where $c$ is a constant, the case (1) is equivalent to case (2), and $t_{4}=t_{3}=0$. ∎
Appendix D OMITTED PROOFS FROM SECTION 6
Proof of Theorem 5
Proof.
Using Lemma 6, we can compute the leader’s utility in the Nash equilibrium, that is $u_{a}^{NE}=\frac{\mu^{*}}{\mu^{*}+1}+\frac{\mu^{*}v_{a1}}{\mu^{*}v_{a1}+v_{b1}}v_{a1}$, where $\mu^{*}$ is a positive solution to $f(\mu)=0$ defined in Lemma 6 and $\mu^{*}\in[z,\frac{v_{b1}z}{v_{a1}}]$. Observe that $u_{a}^{NE}$ is a monotonically increasing function about $\mu^{*}$. Therefore, we have $\frac{x_{a}v_{a1}^{2}}{x_{a}v_{a1}+v_{b1}}+\frac{x_{a}}{x_{a}+1}\leq u_{a}^{NE}\leq\frac{v_{b1}x_{a}v_{a1}}{v_{b1}x_{a}+v_{b1}}+\frac{v_{b1}x_{a}}{v_{b1}x_{a}+v_{a1}}$. Let $\bm{x}_{a}$ denote the leader’s commitment, and $\bm{x}_{b}$ denote the follower’s best response. Then, we can derive that $u_{a}(\bm{x}_{a},\bm{x}_{b})\leq u_{a}(\hat{\bm{x}}_{a},\hat{\bm{x}}_{b})\leq v_{a1}+1$, in which the second inequality sign is obtained by allowing the leader to achieve full utility. Therefore, we have the following inequalities,
$\frac{u_{a}(\bm{x}_{a},\bm{x}_{b})}{\frac{v_{b1}x_{a}v_{a1}}{v_{b1}x_{a}+v_{b1}}+\frac{v_{b1}x_{a}}{v_{b1}x_{a}+v_{a1}}}\leq\frac{u_{a}(\bm{x}_{a},\bm{x}_{b})}{u_{a}^{NE}}\leq\frac{u_{a}(\hat{\bm{x}}_{a},\hat{\bm{x}}_{b})}{u_{a}^{NE}}\leq\frac{v_{a1}+1}{\frac{x_{a}v_{a1}^{2}}{x_{a}v_{a1}+v_{b1}}+\frac{x_{a}}{x_{a}+1}}.$
Consider a certain commitment $\bm{x}_{a}=(\frac{v_{b1}}{v_{b1}+1}x_{a},\frac{1}{v_{b1}+1}x_{a})$, we have $\bm{x}_{b}=(\frac{v_{b1}}{v_{b1}+1},\frac{1}{v_{b1}+1})$. In this commitment, the leader’s utility is
$u_{a}(\bm{x}_{a},\bm{x}_{b})=\frac{\frac{v_{b1}}{v_{b1}+1}x_{a}}{\frac{v_{b1}}{v_{b1}+1}x_{a}+\frac{v_{b1}}{v_{b1}+1}}v_{a1}+\frac{\frac{1}{v_{b1}+1}x_{a}}{\frac{1}{v_{b1}+1}x_{a}+\frac{1}{v_{b1}+1}}=\frac{x_{a}v_{a1}+x_{a}}{x_{a}+1}.$
Substituting $u_{a}(\bm{x}_{a},\bm{x}_{b})$ into the inequality above yields
$\frac{\frac{x_{a}v_{a1}+x_{a}}{x_{a}+1}}{\frac{v_{b1}x_{a}v_{a1}}{v_{b1}x_{a}+v_{b1}}+\frac{v_{b1}x_{a}}{v_{b1}x_{a}+v_{a1}}}\leq\frac{\frac{x_{a}v_{a1}+x_{a}}{x_{a}+1}}{u_{a}^{NE}}\leq\frac{u_{a}(\hat{\bm{x}}_{a},\hat{\bm{x}}_{b})}{u_{a}^{NE}}\leq\frac{v_{a1}+1}{\frac{x_{a}v_{a1}^{2}}{x_{a}v_{a1}+v_{b1}}+\frac{x_{a}}{x_{a}+1}}.$
After simplification, we obtain
$\frac{v_{a1}+1}{v_{a1}+\frac{v_{b1}(x_{a}+1)}{v_{b1}x_{a}+v_{a1}}}\leq\frac{u_{a}(\hat{\bm{x}}_{a},\hat{\bm{x}}_{b})}{u_{a}^{NE}}\leq\frac{v_{a1}+1}{\frac{x_{a}v_{a1}^{2}}{x_{a}v_{a1}+v_{b1}}+\frac{x_{a}}{x_{a}+1}}.$
On the one hand, it is easy to obtain that $\lim_{z\rightarrow\infty}\frac{v_{a1}+1}{\frac{x_{a}v_{a1}^{2}}{x_{a}v_{a1}+v_{b1}}+\frac{x_{a}}{x_{a}+1}}=1$. On the other hand, we let $v_{a1}=x_{a}$ and $v_{b1}=v_{a1}^{2}$, we have $\lim_{x_{a}\rightarrow 0}\frac{v_{a1}+1}{v_{a1}+\frac{v_{b1}(x_{a}+1)}{v_{b1}x_{a}+v_{a1}}}=\infty$. ∎